### v2

# Nonlinear Multi-objective Reinforcement Learning with Provable Guarantees

Abstract
We describe RA-E3 (Reward-Aware Explicit Explore or Exploit), an algorithm with provable guarantees for solving a single or multi-objective Markov Decision Process (MDP) where we want to maximize the expected value of a nonlinear function over accumulated rewards. This allows us to model fairness-aware welfare optimization for multi-objective reinforcement learning as well as risk-aware reinforcement learning with nonlinear Von Neumann-Morgenstern utility functions in the single objective setting. RA-E3 extends the classic E3 algorithm that solves MDPs with scalar rewards and linear preferences. We first state a distinct reward-aware version of value iteration that calculates a non-stationary policy that is approximately optimal for a given model of the environment. This sub-procedure is based on an extended form of Bellman optimality for nonlinear optimization that explicitly considers time and current accumulated reward. We then describe how to use this optimization procedure in a larger algorithm that must simultaneously learn a model of the environment. The algorithm learns an approximately optimal policy in time that depends polynomially on the MDP size, desired approximation, and smoothness of the nonlinear function, and exponentially on the number of objectives.
1 Introduction
Markov Decision Processes (MDPs) model goal-driven interaction with a stochastic environment. The standard problem formulation specifies a scalar-valued reward per time-step and requires an agent to learn a policy (a mapping of states to actions) that maximizes expected (typically time-discounted) total reward. Equivalently, this formulation asks the agent to maximize the expected value of a linear function of total reward. This problem can be solved with provable approximation guarantees in polynomial time with respect to the size of the MDP (Kearns and Singh 2002; Brafman and Tennenholtz 2003; Auer, Jaksch, and Ortner 2008; Azar, Osband, and Munos 2017; Agarwal et al. 2021).
We study the generalization in which we want to optimize a nonlinear function of total reward, where the reward may be vector-valued to represent a multi-objective MDP (MOMDP). We call such a function the welfare function $W$, and we seek to maximize expected welfare $\mathbb{E}\left[W(\mathbf{r})\right]$ where $\mathbf{r}$ is a total reward vector $\mathbf{r}\in\mathbb{R}^{d}$ for a MOMDP with $d$ objectives and the expectation is with respect to randomness in the environment and (possibly) the learned policy. We note that this function is also canonically called the scalarization function within the multiobjective optimization literature (Hayes et al. 2023). We point out at the outset that this nonlinearity is entirely in the objective, as opposed to the many recent works on using nonlinear function approximation as a technique in the optimization of a linear objective over large state spaces.
Nonlinear welfare functions allow us to express a richer space of preferences for the learning agent. For example, in computational social choice, one may express a desire for fairness or balance across objectives by selecting a welfare functions that exhibits diminishing marginal returns such as the Nash Social Welfare $\left(W_{\text{Nash}}(\mathbf{r})=\left(\prod_{i}r_{i}\right)^{1/d}\right)$ (Caragiannis et al. 2019). Even with a single objective, classical decision theory formulates differing risk-attitudes as the maximization of a potentially nonlinear function, also known as a Von Neumann-Morgenstern utility function (Von Neumann and Morgenstern 1947). For example, maximizing the expected value of a concave function of total reward like the logarithm models risk aversion, whereby one prefers a policy that always obtains reward $x$ over a policy that with 50% probability obtains reward $2x$ and otherwise obtains 00. The classic approach of optimizing total scalar reward expresses risk-neutrality, under which the learning agent is indifferent between these two policies.
Contributions and Outline
In this work, we describe the RA-E3 algorithm (Reward-Aware Explicit Explore or Exploit) for solving the problem of maximizing the expected value of a nonlinear function in an MOMDP. To the best of our knowledge, it is the first algorithm to provide provable worst-case approximation guarantees. In the following we summarize our contributions and outline the rest of the paper.
  • In Section 4 we expand the classic model of reinforcement learning to optimize the expected value of a nonlinear welfare function of (possibly) multiple dimensions of reward, motivated by principled multi-objective optimization and modeling of risk attitudes. Our major assumption on the welfare function is that it is smooth. We describe an extended form of Bellman optimality that describes the structure of optimal policies for this problem.
  • In Section 5 we describe Reward-Aware Value Iteration (RA-Value Iteration), a model-based algorithm for computing a non-stationary policy with approximately optimal expected welfare. The algorithm has runtime that is polynomially bounded in the size of the model, horizon time, smoothness of the welfare function, and approximation, and exponential in the number of objectives.
  • In Section 6 we describe how to remove the assumption of a model by incorporating RA-Value Iteration into the explore or exploit framework of the E3 algorithm  (Kearns and Singh 2002). We call the resulting overall algorithm Reward Aware Explicit Explore or Exploit, or RA-E3. In a number of steps bounded polynomially in all relevant parameters except the number of objectives, with high probability, the algorithm will halt in a state $s$ and output a policy $\pi$ that is approximately optimal with respect to the expected welfare beginning at $s$ with time and current reward reset to 0.
2 Related Work
Most algorithms for reinforcement learning focus on the case of a single scalar-valued objective where the goal is to maximize total expected reward (Sutton and Barto 2018). Multi-objective reinforcement learning (Liu, Xu, and Hu 2014) attempts to optimize multiple objectives at once. So-called single-policy methods use a scalarization function to reduce the problem to scalar optimization for a single policy, and we follow this line of research.
The simplest form is linear scalarization, applying a weighted sum on the Q vector (Moffaert, Drugan, and Nowé 2013). A more general problem is to optimize the expected value of a potentially nonlinear function of total reward, which may be vector-valued for a multi-objective optimization context. We refer to such a function as a welfare function following (Barman et al. 2023; Fan et al. 2023; Siddique, Weng, and Zimmer 2020); it is also commonly referred to as a scalarization function as in (Hayes et al. 2023; Agarwal, Aggarwal, and Lan 2022).
Most closely related to our work, (Siddique, Weng, and Zimmer 2020) and (Fan et al. 2023) study algorithms with desirable convergenece properties and empirical performance but no approximation guarantees. (Agarwal, Aggarwal, and Lan 2022) describe another model-based algorithm that can compute approximately optimal policies for a general class of monotone and Lipshitz welfare functions, but rather than maximizing the expected welfare, they maximize the welfare of expected rewards (note the two are not equal for nonlinear welfare functions). Other works have formulated fairness in different ways or settings. For example, (Jabbari et al. 2017) defines an analogue of envy freeness and (Deng et al. 2023) studies a per-time-step fairness guarantee. (Barman et al. 2023) studies welfare maximization in multi-armed bandit problems.
To provide provable approximation and runtime guarantees, we return to the classic E3 algorithm (Kearns and Singh 2002). This result showed that general MDPs could be solved to near-optimality efficiently, meaning in time bounded by a polynomial in the size of the MDP (number of states and actions) and the horizon time. Subsequent results refined the achievable bounds (Brafman and Tennenholtz 2003; Auer, Jaksch, and Ortner 2008; Azar, Osband, and Munos 2017).
3 Preliminaries
A Multi-objective Markov Decision Process (MOMDP) consists of a finite set $\mathcal{S}$ of states, a starting state $s_{1}\in\mathcal{S}$, [^1] In general we may have a distribution over starting states; we assume a single starting state for ease of exposition. a finite set $\mathcal{A}$ of actions, and probabilities $Pr(s^{\prime}|s,a)\in[0,1]$ that determine the probability of transitioning to state $s^{\prime}$ from state $s$ after taking action $a$. Probabilities are normalized so that $\sum_{s^{\prime}}Pr(s^{\prime}|s,a)=1$.
We also have a vector-valued reward function $\mathbf{R}(s,a):\mathcal{S}\times\mathcal{A}\rightarrow[0,1]^{d}$. Each of the $d$ dimensions of the reward vector corresponds to one of the multiple objectives that are to be maximized. At each time step $t$, the agent observes state $s_{t}\in\mathcal{S}$, takes action $a_{t}\in\mathcal{A}(s_{t})$, and receives a reward vector $\bm{R}(s_{t},a_{t})\in\mathbb{R}^{d}$. The environment, in turn, transitions into $s_{t+1}$ with probability $Pr(s_{t+1}|a_{t},s_{t})$.
To make the optimization objective well-posed in MOMDPs with vector-valued rewards, we must specify a scalarization function (Hayes et al. 2023) which we denote as $W:{\mathbb{R}}^{d}\to{\mathbb{R}}$. For fair multi-objective reinforcement learning, we think of each of the $d$ dimensions of the reward vector as corresponding to distinct users. The scalarization function can thus be thought of as a welfare function over the users, and the learning agent is a welfare maximizer. Alternatively, even when $d=1$, a nonlinear function $W$ can be a Von Neumann-Morgenstern utility function (Von Neumann and Morgenstern 1947) that expresses the risk-attitudes of the learning agent, with strictly concave functions expressing risk aversion.
Assumptions
Here we state the preliminary assumptions we make about the reward function $\mathbf{R}(s,a):\mathcal{S}\times\mathcal{A}\rightarrow[0,1]^{d}$ and the welfare function $W:\mathbb{R}^{d}\to\mathbb{R}$.
  1. 1.
    $\mathbf{R}$ is nonnegative and normalized so that the maximum single-step reward in any dimension is 1.
  2. 2.
    $\mathbf{R}$ is deterministic. Note that the environment remains stochastic due to the transition probabilities. For a model-based approach in scalar-valued or linear reinforcement learning, it is sufficient to simply learn the mean of random reward distributions. This is not true when optimizing the expected value of a nonlinear function. We leave full reward distribution learning to future work.
  3. 3.
    $W$ is monotone non-decreasing. That is, for vectors $\mathbf{y}\geq\mathbf{x}$ with $y_{j}\geq x_{j}$ for $1\leq j\leq d$ , $W(\mathbf{y})\geq W(\mathbf{x})$.
  4. 4.
    $W$ is smooth, specifically, uniformly continuous. For all $\epsilon>0$, $\mathbf{x},\mathbf{y}\geq 0$, there exists $\delta_{\epsilon}$ such that
    $|W(\mathbf{x})-W(\mathbf{y})|<\epsilon\quad\text{if}\quad\|\mathbf{x}-\mathbf{y}\|_{1}<\delta_{\epsilon}.$
    Note that uniform continuity is a weaker requirement than Lipshitz continuity.
4 Modeling Optimality
In this section, we expand the classic model of reinforce- ment learning to optimize the expected value of a nonlinear function of (possibly) multiple dimensions of reward. We begin with the notion of a trajectory of state-action pairs.
Definition 1.
Let $M$ be an MOMDP, and let $\pi$ be a policy in $M$. A length $T$ trajectory in $M$ is a tuple $\tau$ of $T$ state-action pairs, that is,
$(s_{1},a_{1}),(s_{2},a_{2}),\dots,(s_{T},a_{T}).$
For $1\leq k\leq k^{\prime}\leq T$, let $\tau_{k:k^{\prime}}$ be the sub-trajectory consisting of pairs $(s_{k},a_{k}),\dots,(s_{k^{\prime}},a_{k^{\prime}})$. Let $\tau_{0;0}$ denote the empty trajectory.
For a discount factor $\gamma$, we calculate the total discounted reward of a trajectory. Note that this is a vector in general.
Definition 2.
For length $T$ trajectory $\tau$ and discount factor $10\leq\gamma<$ 1, the total discounted reward along $\tau$ is the the vector
$\bm{R}(\tau,\gamma)=\sum_{t=1}^{T}\gamma^{t-1}\bm{R}(s_{t},a_{t}).$
For ease of exposition we will frequently leave $\gamma$ implicit from context and simply write $\bm{R}(\tau)$.
A policy is a function $\pi(a\mid\tau,s)\in[0,1]$ mapping past trajectories and current states to probability distributions over actions, that is, $\sum_{a}\pi(a\mid\tau,s)=1$ for all $\tau$ and $s$. A stationary policy is the special case of a policy that depends only on the current state: $\pi(a\mid s)$.
Definition 3.
The probability that a $T$-trajectory $\tau$ is traversed in an MOMDP $M$ upon starting in state $s_{1}$ and executing policy $\pi$ is
$Pr^{\pi}[\tau]$ $=\pi(a_{1}\mid\tau_{0:0},s_{1})\times$
$\qquad\prod_{t=2}^{T}\pi\left(a_{t}~{}\bigg{|}~{}\tau_{1:t-1},s_{t}\right)Pr(s_{t}|s_{t-1},a_{t-1}).$
Problem Formulation
Given a policy, a finite time-horizon $T$, and a starting state $s_{1}$ we can calculate the expected welfare of total discounted reward along a trajectory as follows. Our goal is to maximize this quantity. That is, we want to compute a policy that maximizes the expected $T$-step discounted welfare.
Definition 4.
For a policy $\pi$ and a start state $s$, the expected $T$-step discounted welfare is
$\mathbb{E}_{\tau\sim\pi}\bigl{[}W\bigl{(}\bm{R}(\tau)\bigr{)}\bigr{]}=\sum_{\tau}Pr^{\pi}[\tau]W(\mathbf{R}(\tau))$
where the expectation is taken over all length $T$ trajectories beginning at $s$.
Note that this objective is not equal to $W\left(\mathbb{E}_{\tau\sim\pi}\bigl{[}\bm{R}(\tau)\bigr{]}\right)$, which others have studied (Agarwal, Aggarwal, and Lan 2022; Siddique, Weng, and Zimmer 2020), for a nonlinear $W$. The former (our objective) is also known as expected scalarized return (ESR) whereas the latter is also known as scalarized expected return (SER) (Hayes et al. 2023). While SER makes sense for a repeated decision-making problem, it does not optimize for expected welfare for any particular trajectory, in contrast to ESR. To the best of our knowledge, our work constitutes the first algorithm with worst-case approximation and polynomial runtime guarantees for ESR.
Form of Optimal Policy and Value Functions
The optimal policy for this finite time horizon setting is a function also of the number of time steps remaining. We write such a policy as $\pi(a\mid s,\tau,t)$ where $\tau$ is a trajectory (the history), $s$ is the current state, and $t$ is the number of time steps remaining, $i.e.$ $t=T-|\tau|$.
We can similarly write the extended value function of a policy $\pi$ in MORL context. We write $\tau$ as the history or trajectory prior to some current state $s$ and $\tau^{\prime}$ as the future, the remaining $t$ steps determined by the policy $\pi$ and the environmental transitions.
Definition 5.
The value of a policy $\pi$ beginning at state $s$ after history $\tau$ and executing $t$ more actions is
$V^{\pi}(s,\tau,t)=\mathbb{E}_{\tau^{\prime}\sim\pi}\left[W(\mathbf{R}(\tau)+\gamma^{T-t}\mathbf{R}(\tau^{\prime}))\right]$
where the expectation is taken over all length $t$ trajectories $\tau^{\prime}$ beginning at $s$. The optimal value function is
$V^{*}(s,\tau,t)=\max_{\pi}V^{\pi}(s,\tau,t)$
and the optimal policy is
$\pi^{*}\in\operatorname*{arg\,max}_{\pi}V^{\pi}(s,\tau,t).$
Note that because $W$ is nonlinear, the value function $V^{\pi}$ cannot be decomposed in the same way as in the traditional Bellman equations. Instead, we develop an extended form of Bellman optimality by showing that the optimal value function can be written as a function of current state $s$, accumulated discounted reward $\mathbf{R}(\tau)$, and number of timesteps remaining in the task $t$.
Lemma 6.
Let $\mathcal{V}(s,\mathbf{R}(\tau),0)=W(\mathbf{R}(\tau))$ for all states $s$ and trajectories $\tau$. For every state $s$, history $\tau$, and $t>0$ time steps remaining, let
$\mathcal{V}(s,\mathbf{R}(\tau),t)=\max_{a}\mathbb{E}_{s^{\prime}}\left[\mathcal{V}(s^{\prime},\mathbf{R}(\tau)+\gamma^{T-t}\mathbf{R}(s,a),t-1)\right].$
Then $V^{*}(s,\tau,t)=\mathcal{V}(s,\mathbf{R}(\tau),t)$.
Proof.
We proceed by induction on $t$. In the base case of $t=0$, we have simply that $\mathcal{V}(s,\mathbf{R}(\tau),0)=W(\mathbf{R}(\tau))$ by the definition of $\mathcal{V}$. But when $t=0$, the agent is at the terminal state, which implies any $\tau^{\prime}$ in Definition 5 will be the empty trajectory $\tau_{0;0}$. So
$V^{*}(s,\tau,0)=W(\mathbf{R}(\tau)+\gamma^{T-t}\mathbf{R}(\tau_{0:0}))=W(\mathbf{R}(\tau)).$
For the inductive hypothesis, suppose that the equality holds for all states $s$ and histories $\tau$ for up to $t-1\geq 0$ time steps remaining, i.e.
$V^{*}(s,\tau,t-1)=\max_{a}\mathbb{E}_{s^{\prime}}\left[\mathcal{V}(s^{\prime},\mathbf{R}(\tau)+\gamma^{T-t}\mathbf{R}(s,a),t-2)\right].$
Then with $t$ steps remaining:
$V^{*}(s,\tau,t)=\max_{\pi}\sum_{\tau^{\prime}}Pr^{\pi}(\tau^{\prime})W(\bm{R}(\tau)+\gamma^{T-t}\bm{R}(\tau^{\prime}))$
where the sum is taken over all length-$t$ trajectories $\tau^{\prime}$ beginning at $s$, and $Pr^{\pi}(\tau^{\prime})$ is the probability that $\tau^{\prime}$ is traversed under policy $\pi$. Note that we can further decompose $\tau^{\prime}$ into
$\tau^{\prime}=(s,a)\oplus\tau^{\prime\prime}$
where $(s,a)$ denotes the state-action pair corresponding to the current timestep, $\oplus$ concatenates, and $\tau^{\prime\prime}$ denotes a trajectory of length $t-1$ beginning at some state $s^{\prime}$ transitioned from state $s$ via taking action $a$. Then we can rewrite $V^{*}$ as
$V^{*}(s,\tau,t)=\max_{\pi}\sum_{a}$ $\pi(a\mid s,\tau,t)~{}\mathbb{E}_{s^{\prime}}\bigg{[}\sum_{\tau^{\prime\prime}}Pr^{\pi}(\tau^{\prime\prime})W(\bm{R}(\tau)$
$+\gamma^{T-t}\bm{R}(s,a)+\gamma^{T-t+1}\bm{R}(\tau^{\prime\prime}))\bigg{]}$
where the latter sum is taken over all length $t-1$ trajectories $\tau^{\prime\prime}$ beginning at $s^{\prime}$ and the expectation is taken over the environmental transitions to $s^{\prime}$ from $s$ given $a$.
Note that for all actions $a$, $\pi(a\mid s,\tau,t)\in[0,1]$ can be chosen independently of any terms that appears in $Pr^{\pi}(\tau^{\prime\prime})$ for any $\tau^{\prime\prime}$ by the decomposition of $\tau^{\prime}$ above. This implies that at the current timestep, the maximizer (or the optimal policy) should choose action $a$ that maximizes the quantity
$\mathbb{E}_{s^{\prime}}\bigg{[}\sum_{\tau^{\prime\prime}}Pr^{\pi}(\tau^{\prime\prime})W(\bm{R}(\tau)+\gamma^{T-t}\bm{R}(s,a)+\gamma^{T-t+1}\bm{R}(\tau^{\prime\prime}))\bigg{]}$
with probability 1. Therefore, the expression for $V^{*}(s,\tau,t)$ can be rewritten as
$V^{*}(s,\tau,t)$ $=\max_{a}\mathbb{E}_{s^{\prime}}\bigg{[}\max_{\pi}\sum_{\tau^{\prime\prime}}Pr^{\pi}(\tau^{\prime\prime})W(\bm{R}(\tau)$
$\qquad+\gamma^{T-t}\bm{R}(s,a)+\gamma^{T-(t-1)}\bm{R}(\tau^{\prime\prime}))\bigg{]}$
$=\max_{a}\mathbb{E}_{s^{\prime}}\left[V^{*}(s^{\prime},\tau\oplus(s,a),t-1)\right]$
$=\max_{a}\mathbb{E}_{s^{\prime}}\left[\mathcal{V}(s^{\prime},\bm{R}(\tau)+\gamma^{T-t}\bm{R}(s,a),t-1)\right]$
$=\mathcal{V}(s,\bm{R}(\tau),t).$
where the second equality comes from the definition of $V^{*}$, the third equality comes from the inductive hypothesis, and the fourth equality comes from the definition of $\mathcal{V}$. ∎
By Lemma 6, we can parameterize the optimal value function for MORL by the current state $s$, accumulated reward $\mathbf{R_{acc}}\in{\mathbb{R}}^{d}$, and number of timesteps remaining $t$. We will use this formulation of $V^{*}$ in the remainder of the paper.
Definition 7 (Recursive Formulation of $V^{*}$).
Let $\mathbf{R_{acc}}=\mathbf{R}(\tau)$ be the vector of accumulated reward along a history prior to some state $s$. Let $V^{*}(s,\mathbf{R_{acc}},0)=W(\mathbf{R_{acc}})$ for all $s$ and $\mathbf{R_{acc}}$. For $t>0$,
$V^{*}(s,\mathbf{R}_{acc},t)$ $=\max_{a}\sum_{s^{\prime}}Pr(s^{\prime}|s,a)\times$
$\qquad V^{*}(s^{\prime},\mathbf{R}_{acc}+\gamma^{T-t}\mathbf{R}(s,a),t-1).$
Before proceeding we want to provide some intuition as to why the traditional form of Bellman optimality conditions do not hold. The same reasoning helps to explain why stationary policies are not, in general, optimal for our objective.
Suppose we are optimizing the product of the reward between two objectives (i.e., the welfare function is the product or geometric mean), and at some state $s$ with some prior history $\tau$ we can choose between two policies $\pi_{1}$ or $\pi_{2}$. Suppose that the future discounted reward vector under $\pi_{1}$ is $(1,1)$, whereas it is $(10,0)$ under $\pi_{2}$. So $\pi_{1}$ has greater expected future welfare and traditional Bellman optimality would suggest we should choose $\pi_{1}$. However, if $\mathbf{R}(\tau)=(0,10)$, we would actually be better off in terms of total welfare choosing $\pi_{2}$. In other words, both past and future reward are relevant when optimizing for the expected value of a nonlinear welfare function.
Horizon Time
Note that an approximation algorithm for this discounted finite time horizon problem can also be used as an approximation algorithm for the discounted infinite time horizon problem. Informally, discounting by $\gamma$ with bounded maximum reward implies that the first $T\approx 1/(1-\gamma)$ steps dominate overall returns.
Lemma 8.
Let $M$ be any MOMDP, and let $\pi$ be any policy in $M$. Assume the welfare function $W:\mathbb{R}^{d}\to\mathbb{R}$ is uniformly continuous and $\mathbf{R}(s,a)\in[0,1]^{d}~{}\forall s,a$. Then for all $\epsilon>0$, there exists $\delta_{\epsilon}>0$ such that if
$T\geq\left(\frac{1}{1-\gamma}\right)\log\left(\frac{1}{\delta_{\epsilon}(1-\gamma)}\right)$
then for any state $s,$
$V^{\pi}(s,\mathbf{0},T)\leq\lim_{T\to\infty}V^{\pi}(s,\mathbf{0},T)\leq V^{\pi}(s,\mathbf{0},T)+\epsilon.$
We call the value of the lower bound on $T$ given above the $\epsilon$-horizon time for the MOMDP $M$.
Proof.
[…]
$\bigg{\|}\sum_{k=1}^{\infty}\gamma^{k-1}\bm{R}(s_{k},a_{k})-\sum_{k=1}^{T}\gamma^{k-1}\bm{R}(s_{k},a_{k})\bigg{\|}$
$=$ $\bigg{\|}\sum_{k=T+1}^{\infty}\gamma^{k-1}\bm{R}(s_{k},a_{k})\bigg{\|}$
$=$ $\gamma^{T}\bigg{\|}\sum_{k=1}^{\infty}\gamma^{k-1}\bm{R}(s_{T+k},a_{T+k})\bigg{\|}$
$\leq$ $\gamma^{T}\left(\frac{1}{1-\gamma}\right)$
Solving
$\gamma^{T}\left(\frac{1}{1-\gamma}\right)\leq\delta_{\epsilon}$
for $T$ yields the desired bound on $T$. Since the inequality holds for every fixed trajectory, it also holds for the distribution over trajectories induced by any policy $\pi$. ∎
5 The RA-Value Iteration Algorithm
In this section, we describe a model-based algorithm called RA-Value Iteration that efficiently finds an approximately optimal policy. Note that the values of $V^{*}$ with $T$ timesteps remaining only depends on the those with $T-1$ timesteps remaining. This naturally induces a dynamic programming approach. A naive algorithm for learning a nonstationary policy would need to consider all possible prior trajectories for each decision point, leading to a runtime complexity that depends on the exponential term $|\mathcal{S}|^{T}$.
Instead, our algorithm is derived from the recursive definition of the optimal multi-objective value function $V^{*}$ in 7 justified by Lemma 6, parameterized by the accumulated discounted reward vector $\mathbf{R_{acc}}$ instead of the prior history. Note that even if rewards were integer, $\mathbf{R_{acc}}$ would not be due to discounting. We must therefore introduce a discretization method to our dynamic programming algorithm that treats “nearby” accumulated reward vectors as equivalent. It is parameterized by some $\alpha\in(0,1)$ where a smaller $\alpha$ leads to a finer discretization but increases the runtime.
Definition 9.
For a given discretization precision parameter $\alpha\in(0,1)$, define $f_{\alpha}:{\mathbb{R}}^{d}\to(\alpha\mathbb{Z})^{d}$ by
$f_{\alpha}(\mathbf{R})=\left(\left\lfloor\frac{R_{1}}{\alpha}\right\rfloor\cdot\alpha,\left\lfloor\frac{R_{2}}{\alpha}\right\rfloor\cdot\alpha,\cdots,\left\lfloor\frac{R_{d}}{\alpha}\right\rfloor\cdot\alpha\right).$
In other words, $f$ maps any $d$-dimension vector to its nearest $\alpha$ multiples. We are now prepared to define the algorithm.
Algorithm 1 The RA-Value Iteration Algorithm
1:Parameters: Discretization precision $\alpha\in(0,1)$, Discount factor $\gamma\in[0,1)$, Reward dimension $d$, finite time horizon $T$, uniformly continuous welfare function $W$, discretization function $f$.
2:Require: Normalize $\mathbf{R}(s,a)\in[0,1]^{d}$ for all $s\in\mathcal{S},a\in\mathcal{A}$. Initialize $V(s,\mathbf{R_{acc}},0)=W(\mathbf{R_{acc}})$ for all $\mathbf{R_{acc}}\in\{0,\mathbf{\alpha},2\mathbf{\alpha},\dots,\lfloor T/\alpha\rfloor\cdot\alpha\}^{d}$, $s\in\mathcal{S}$.
3:for t = 1 to $T$ do
4:     for $\mathbf{R_{acc}}\in\{0,\mathbf{\alpha},2\mathbf{\alpha},\dots,\lfloor(T-t)/\alpha\rfloor\cdot\alpha\}^{d}$ do
5:         for all $s\in\mathcal{S}$ do
6:
$V(s,\mathbf{R_{acc}},t)$ $\leftarrow\max_{a}\sum_{s^{\prime}}Pr(s^{\prime}|s,a)\times$
$\quad V(s^{\prime},f_{\alpha}(\mathbf{R}_{acc}+\gamma^{T-t}\mathbf{R}(s,a)),t-1)$
7:
$\pi(s,\mathbf{R_{acc}},t)$ $\leftarrow\operatorname*{arg\,max}_{a}\sum_{s^{\prime}}Pr(s^{\prime}|s,a)\times$
$\quad V(s^{\prime},f_{\alpha}(\mathbf{R}_{acc}+\gamma^{T-t}\mathbf{R}(s,a)),t-1)$
8:         end for
9:     end for
10:end for
The runtime of the RA-Value Iteration algorithm is $O\left(|\mathcal{S}||\mathcal{A}|(T/\alpha)^{d}\right)$ which is polynomial time for constant dimensionality $d$; recall that $d$ is the number of objectives, possibly just 1.
We now prove a worst-case approximation guarantee for the RA-Value Iteration algorithm: given a smooth welfare function $W$, the algorithm computes a non-stationary policy with approximately optimal expected welfare. We begin by proving an important structural property: the optimal value function inherits smoothness of the welfare function, i.e. the optimal value function will be smooth if the welfare function is smooth.
Lemma 10 (Uniform continuity of multi-objective value function).
Let the welfare function $W:\mathbb{R}^{d}\to\mathbb{R}$ be uniformly continuous. Fix $s\in\mathcal{S}$ and $t\in\{0,1,\dots,T\}$, then for all $\epsilon>0$, there exists $\delta_{\epsilon}>0$ such that
[…]
Proof.
Without loss of generality, assume $V^{*}(s,\mathbf{R_{1}},t)\geq V^{*}(s,\mathbf{R_{2}},t)$. Let $\pi_{1}$ be the induced optimal policy by $V^{*}(s,\mathbf{R_{1}},t)$, i.e.
$V^{*}(s,\mathbf{R_{1}},t)=V^{\pi_{1}}(s,\mathbf{R_{1}},t).$
Then by the optimality of $V^{*}$,
$V^{\pi_{1}}(s,\mathbf{R_{2}},t)$ $\leq V^{*}(s,\mathbf{R_{2}},t)$
$=\max_{\pi}\sum_{\tau}Pr^{\pi}[\tau]W(\mathbf{R_{2}}+\mathbf{R}(\tau)).$
Since $W$ is uniformly continuous, there exists $\delta_{\epsilon}>0$ such that $|W(\mathbf{x})-W(\mathbf{y})|<\epsilon$ if $\|\mathbf{x}-\mathbf{y}\|<\delta_{\epsilon}$. Then
$\bigg{|}V^{*}(s,\mathbf{R_{1}},t)-V^{*}(s,\mathbf{R_{2}},t)\bigg{|}$
$\leq$ $\bigg{|}V^{\pi_{1}}(s,\mathbf{R_{1}},t)-V^{\pi_{1}}(s,\mathbf{R_{2}},t)\bigg{|}$
$=$ $\bigg{|}\sum_{\tau}Pr^{\pi_{1}}[\tau]W(\mathbf{R_{1}}+\mathbf{R}(\tau))-\sum_{\tau}Pr^{\pi_{1}}[\tau]W(\mathbf{R_{2}}+\mathbf{R}(\tau))\bigg{|}$
$=$ $\sum_{\tau}Pr^{\pi_{1}}[\tau]\bigg{|}W(\mathbf{R_{1}}+\mathbf{R}(\tau))-W(\mathbf{R_{2}}+\mathbf{R}(\tau))\bigg{|}$
where the sum is over all $t$-trajectories $\tau$ that start in state $s$. If $\|\mathbf{R_{1}}-\mathbf{R_{2}}\|<\delta_{\epsilon}$ then we have
$\bigg{|}W(\mathbf{R_{1}}+\mathbf{R}(\tau))-W(\mathbf{R_{2}}+\mathbf{R}(\tau))\bigg{|}<\epsilon.$
Therefore,
$\bigg{|}V^{*}(s,\mathbf{R_{1}},t)-V^{*}(s,\mathbf{R_{2}},t)\bigg{|}<\epsilon$
since the sum of probabilities over all $t$-trajectories $\tau$ in $M$ that start in state $s$ induced by $\pi_{1}$ must add up to $1$. ∎
We now prove the approximation guarantee of RA-Value Iteration, that it achieves an additive error that scales with the smoothness of $W$ and the number of time steps remaining.
Lemma 11 (Approximation Error of RA-Value Iteration).
For uniformly continuous welfare function $W$, for all $\epsilon>0$, there exists $\alpha_{\epsilon}$ such that
$V(s,\mathbf{R_{acc}},t)\geq V^{*}(s,\mathbf{R_{acc}},t)-t\epsilon$
$\forall s\in\mathcal{S},\mathbf{R_{acc}}\in{\mathbb{R}}^{d},t\in\{0,1,\dots,T\}$, where $V(s,\mathbf{R_{acc}},t)$ is the value function calculated by Algorithm 1.
Proof.
We prove this additive error bound by induction on the number of timesteps $t$ remaining in the task.
[…]
$=V^{*}(s,\mathbf{R_{acc}},0)\quad\forall s,\mathbf{R_{acc}}$
Inductive hypothesis: for $t=0,\dots,k-1$,
$V(s,\mathbf{R_{acc}},t)\geq V^{*}(s,\mathbf{R_{acc}},t)-t\epsilon$
Then consider with $k$ steps remaining.
$V(s,\mathbf{R_{acc}},k)$
$=$ $\max_{a}\sum_{s^{\prime}}Pr(s^{\prime}|s,a)V(s^{\prime},f_{\alpha}(\mathbf{R_{acc}}+\gamma^{T-k}\mathbf{R}(s,a)),k-1)$
$\geq$ $\max_{a}\sum_{s^{\prime}}Pr(s^{\prime}|s,a)\times$
$\bigg{(}V^{*}(s^{\prime},f_{\alpha}(\mathbf{R_{acc}}+\gamma^{T-k}\mathbf{R}(s,a)),k-1)-(k-1)\epsilon\bigg{)}.$
where the last line follows from the inductive hypothesis. Since $W$ is uniformly continuous by Lemma 10, there exists $\delta_{\epsilon}$ such that
$\bigg{|}V^{*}\bigg{(}s^{\prime},f_{\alpha}(\mathbf{R_{acc}}$ $+\gamma^{T-k}\mathbf{R}(s,a)),k-1\bigg{)}$
$-V^{*}\bigg{(}s^{\prime},\mathbf{R_{acc}}+\gamma^{T-k}\mathbf{R}(s,a),k-1\bigg{)}\bigg{|}<\epsilon$
if
$\bigg{\|}f_{\alpha}(\mathbf{R_{acc}}+\gamma^{T-k}\mathbf{R}(s,a))-(\mathbf{R_{acc}}+\gamma^{T-k}\mathbf{R}(s,a))\bigg{\|}<\delta_{\epsilon}.$
Hence, it suffices to choose $\alpha=\delta_{\epsilon}/d$, which implies
$\bigg{\|}f_{\alpha}(\mathbf{R_{acc}}+\gamma^{T-k}\mathbf{R}(s,a))-(\mathbf{R_{acc}}$ $+\gamma^{T-k}\mathbf{R}(s,a))\bigg{\|}$
$<(\delta_{\epsilon}/d)\cdot d<\delta_{\epsilon}.$
Then
$V(s,\mathbf{R_{acc}},k)$
$\geq$ $\max_{a}\sum_{s^{\prime}}Pr(s^{\prime}|s,a)V^{*}\bigg{(}s^{\prime},\mathbf{R_{acc}}+\gamma^{T-k}\mathbf{R}(s,a),k-1\bigg{)}$
$\quad-\epsilon-(k-1)\epsilon$
$=$ $V^{*}(s,\mathbf{R}_{acc},k)-k\epsilon.$
Theorem 12 (Optimality Guarantee of RA-Value Iteration).
For a uniformly continuous welfare function $W$, there exists an algorithm $A$, taking inputs $\epsilon$, $|\mathcal{S}|$, and $|\mathcal{A}|$, such that the total number of actions and computation time taken by $A$ is polynomial in $1/\epsilon$, $|\mathcal{S}|$, $|\mathcal{A}|$, the horizon time $T=1/(1-\gamma)$, and exponential in the number of objectives $d$, and $A$ will output a policy $\hat{\pi}$, such that $V^{\hat{\pi}}(s,0,T)\geq V^{*}(s,0,T)-\epsilon$.
Proof.
Note that in particular, Lemma 11 implies that $V(s,\mathbf{0},T)\geq V^{*}(s,\mathbf{0},T)-T\epsilon$. This implies that for a finite time horizon $T$ and a desired approximation error $\epsilon$, it suffices to choose $\alpha_{T\epsilon}=\delta_{T\epsilon}/d$ where $\delta_{T\epsilon}$ comes from the smoothness of the welfare function. One then uses RA-Value Iteration. ∎
6 Removing Knowledge of the Model
We have shown that RA-value Iteration can find an approximately optimal policy in polynomial time (for a constant number of objectives) given access to a model of the environment. Next, we describe how to learn the model by extending the classical E3 algorithm (Kearns and Singh 2002) to the multi-objective case. We call the resulting algorithm Reward-Aware Explicit Explore or Exploit (RA-E3).
A learning algorithm cannot recover from low return during the early part of learning due to the exponentially decaying effects of the discounting factor. Therefore, our goal for a learning algorithm is to efficiently compute a policy while exploring so that, starting from the current state with 0 initial accumulated reward and time steps remaining reset to $T$, that policy has expected welfare within $\epsilon$ of the best possible beginning from that state.
For ease of exposition, we will describe the algorithm under the assumption that we know the optimal return $V^{*}(s,\mathbf{0},T)$. This assumption can be further removed via an exploration-biased approach during the offline policy computation stage similar to the original E3 algorithm (Kearns and Singh 2002). We can now state the main result yielded by the RA-E3 algorithm.
Theorem 13 (RA-E3).
Let $V^{*}(s,0,T)$ denote the value function for the policy with the optimal expected welfare in the MOMDP $M$ starting at state $s$, with $\mathbf{0}\in{\mathbb{R}}^{d}$ accumulated reward and $T$ timesteps remaining. Then for a uniformly continuous welfare function $W$, there exists an algorithm $A$, taking inputs $\epsilon$, $\beta$, $|\mathcal{S}|$, $|\mathcal{A}|$, and $V^{*}(s,\mathbf{0},T)$, such that the total number of actions and computation time taken by $A$ is polynomial in $1/\epsilon$, $1/\beta$, $|\mathcal{S}|$, $|\mathcal{A}|$, the horizon time $T=1/(1-\gamma)$ and exponential in the number of objectives $d$, and with probability at least $1-\beta$, $A$ will halt in a state $s$, and output a policy $\hat{\pi}$, such that $V^{\hat{\pi}}_{M}(s,0,T)\geq V^{*}(s,0,T)-\epsilon$.
The algorithm is mostly a straightforward extension of the E3 algorithm expanded to incorporate a multi-objective (state, action)-pair value function. Most proofs are identical apart from this or the addition of extra $|\mathcal{A}|$ runtime factors. The notable exception are the exploitation subroutines which we have introduced already. Here we sketch the argument by outlining the entire RA-E3 algorithm and emphasizing the role of our RA-Value Iteration optimization algorithm. We provide additional details comparing our analysis with that of (Kearns and Singh 2002) in the appendix.
Introducing RA-E3
The extension works as follows. The algorithm starts off by doing balanced wandering (Kearns and Singh 2002): when encountering a new state, the algorithm selects a random action. However, when revisiting a previously visited state, it chooses the least attempted action from that state, resolving ties by random action selection. At each state-action pair $(s,a)$ it tries, the algorithm stores the reward $\mathbf{R}(s,a)$ received and an estimate of the transition probabilities $Pr(s^{\prime}\mid s,a)$ derived from the empirical distribution of next states reached during balanced wandering.
Next, we introduce the notion of a known state (Kearns and Singh 2002), which refers to a state that the algorithm has explored to the extent that the estimated transition probabilities for any action from that state closely approximate their actual values. Denote the number of times a state needs to be visited as $m_{known}$. We will specify the value later in our runtime characterization.
States are thus categorized into three groups: known states, which the algorithm has extensively visited and obtained reliable transition statistics; states that have been visited before but remain unknown due to limited trials and therefore unreliable data; and states that have not been explored at all. By the Pigeonhole Principle, accurate statistics will eventually accumulate in some states over time, leading to their becoming known. Let $S$ be the set of currently known states, the algorithm can build the current known-state MOMDP $M_{S}$ that is naturally induced on $S$ by the full MOMDP $M$ with all “unknown” states merged into a single absorbing state. Although the algorithm cannot access $M_{S}$ directly, it will have an approximation $\hat{M}_{S}$ by the definition of the known states. By the simulation lemma (Kearns and Singh 2002), $\hat{M}_{S}$ will be an accurate model in the sense that the expected $T$-step welfare of any policy in $\hat{M}_{S}$ is close to its expected $T$-step return in $M_{S}$. (Here $T$ is the horizon time.) Hence, at any timestep, $\hat{M}_{S}$ functions as an incomplete representation of $M$, specifically focusing on the part of $M$ that the algorithm possesses a strong understanding of.
This is where we insert RA-Value Iteration. The algorithm performs the two off-line optimal policy computations; i) first on $\hat{M}_{S}$ using RA-Value Iteration to compute an exploitation policy that yields an approximately optimal welfare and ii) second performing traditional value iteration on $\hat{M}^{\prime}_{S}$, which has the same transition probabilities as $\hat{M}_{S}$, but different payoffs: in $\hat{M}^{\prime}_{S}$, the absorbing state (representing “unknown” states) has scalar reward $1$ and all other states have scalar reward 00. The optimal policy in $\hat{M}^{\prime}_{S}$ simply exits the known model as rapidly as possible, rewarding exploration.
By the explore or exploit lemma (Kearns and Singh 2002), the algorithm is guaranteed to either output a policy with approximately optimal return in $M$, or to improve the statistics at an unknown state. Again by the Pigeonhole Principle, a new state becomes known after the latter case occurs for some finite number of times, and thus the algorithm is always making progress. In the worst case, the algorithm builds a model of the entire MOMDP $M$. Having described the elements, we now outline the entire extended algorithm, where the notations are consistent with (Kearns and Singh 2002).
RA-E3 Algorithm:
  • (Initialization) Initially, the set $S$ of known states is empty.
  • (Balanced Wandering) Any time the current state is not in $S$, the algorithm performs balanced wandering
  • (Discovery of New Known States) Any time a state $s$ has been visited $m_{known}$ (Kearns and Singh 2002) times during balanced wandering, it enters the known set $S$, and no longer participates in balanced wandering.
  • (Off-line Optimizations) Upon reaching a known state $s\in S$ during balanced wandering, the algorithm performs the two off-line optimal policy computations on $\hat{M}_{S}$ and $\hat{M}^{\prime}_{S}$ described above:
    • (Attempted Exploitation) Use RA-Value Iteration algorithm to compute an $\epsilon/2$-optimal policy on $\hat{M}_{S}$. If the resulting exploitation policy $\hat{\pi}$ achieves return from $s$ in $\hat{M}_{S}$ that is at least $V^{*}(s,\mathbf{0},T)-\epsilon/2$, the algorithm halts and outputs $\hat{\pi}$.
    • (Attempted Exploration) Otherwise, the algorithm executes the resulting exploration policy derived from the off-line computation on $\hat{M}^{\prime}_{S}$ for $T$ steps in $M$.
  • (Balanced Wandering) Any time an attempted exploitation or attempted exploration visits a state not in $S$, the algorithm resumes balanced wandering.
This concludes the description of the algorithm. The failure parameter $\beta$ in the statement of the main theorem can be handled similar to the original algorithm, giving a maximum number of attempted explorations of
$O((G^{T}_{\max}/\epsilon)|\mathcal{S}|\log(|\mathcal{S}|/\beta)m_{known}),$
where $G^{T}_{\max}\leq W(\mathbf{T})$ denotes the maximum possible $T$-step welfare (recall $\mathbf{R}(s,a)\in[0,1]^{d}~{}\forall s,a$).
$m_{known}=O(((|\mathcal{S}||\mathcal{A}|TG^{T}_{\max})/\epsilon)^{4}|\mathcal{A}|\log(1/\beta))$ by the same analysis as (Kearns and Singh 2002). Note that this bound differs from the original one by a factor of $|\mathcal{A}|^{5}$. This is because i) we no longer assume the number of actions be a constant as in the original assumption, which accounts for $|\mathcal{A}|$ outside the quartic, and ii) the optimal policy $\pi$ is no longer stationary with respect to the current state, which accounts for the $|\mathcal{A}|$ inside the quartic.
Thus, the total computation time is bounded by $O\left(|\mathcal{S}||\mathcal{A}|(T/\alpha_{T\epsilon/2})^{d}\right)$ (the time required for RA-Value Iteration during off-line computations with $\epsilon/2$ precision by Lemma 11) times $T$ times the maximum number of attempted explorations, giving
$O((|\mathcal{S}|^{2}|\mathcal{A}|TG^{T}_{\max}/\epsilon)(T/\alpha_{T\epsilon/2})^{d}\log(|\mathcal{S}|/\beta)m_{known}).$
This concludes the sketch of the RA-E3 theorem.
Appendix
Appendix A Example and Practical Remarks
The RA-E3 algorithm has many parameters and terms in the analysis. In this section, we provide a concrete example of a particular welfare function, the setting of relevant parameters, and a derivation of a simplified runtime analysis.
Smoothed Proportional Fairness Welfare Function.
We will use the smoothed proportional fairness objective for our example: $W_{SPF}(\mathbf{x})=\sum_{i=1}^{d}\ln(x_{i}+1)$. The sum of logarithms is known as the proportional fairness objective - note that it is simply the log-transform of the Nash Social Welfare (or geometric mean) with better numerical stability. The addition of 1 inside the log is a smoothing factor intended to stabilize behavior near 0.
Calculating $\delta_{\epsilon}$.
Uniform continuity is a weaker assumption than the more common form of $L$-Lipshitz continuity for some constant $L$. Note that for an $L$-Lipshitz function (equivalently, for a differentiable function with the magnitude of the gradient bounded everywhere on the domain by $L$), then the correct value of $\delta_{\epsilon}$ is just $\epsilon/L$, where $\epsilon$ is the desired approximation factor of the algorithm.
By taking the gradient, we can see that the $W_{SPF}(\mathbf{x})$ is Lipshitz-continuous with Lipshitz constant $d$ for nonnegative $\mathbf{x}$, where $d$ is the dimension of vector reward, so $\delta_{\epsilon}=\epsilon/d$.
Why The Discount Factor $\gamma$ Matters: Calculating The Horizon Time $T$ and maximum return $G^{T}_{max}$.
For tasks with a finite time Horizon, $T$ is a modeling choice. Otherwise, for a discounting factor $\gamma<1$, $T$ must be large enough to ensure that planning beyond the horizon time is inconsequential in terms of total discounted reward. By Lemma 8, we need
$T\geq\left(\frac{1}{1-\gamma}\right)\log\left(\frac{1}{\delta_{\epsilon}(1-\gamma)}\right)$
which for our running example, can be simplified to
$T\geq\left(\frac{1}{1-\gamma}\right)\log\left(\frac{d}{\epsilon(1-\gamma)}\right).$
Recall that $G^{T}_{max}$ is the maximum $T$-step welfare possible, trivially bounded by the welfare of $T$ times the identity vector (recall our rewards are normalized to at most 1 in every dimension). So in our running example,
$G^{T}_{max}\leq d\ln(1+T).$
Setting The Discretization Parameter $\alpha_{T\epsilon}$ in The RA-Value Iteration Algorithm.
Recall from Theorem 12 that we should set the discretization parameter in the RA-Value Iteration Algorithm to $\alpha_{T\epsilon}=\delta_{T\epsilon}/d$ where $\delta_{T\epsilon}$ is sufficient to drive the Lemma 11 bound to $\epsilon$, that is, it should be
$\alpha_{T\epsilon}=\frac{\delta_{\epsilon}}{T\cdot d}.$
In our running example, since $\delta_{\epsilon}=\frac{\epsilon}{d}$ so we should set
$\alpha_{T\epsilon}=\frac{\epsilon}{T\cdot d^{2}}.$
Note that concave welfare functions exhibit better smoothness by moving further away from the origin. A further optimization of the RA-Value Iteration algorithm can exploit this property by “coarsening” the grid further away from the origin, i.e. reducing grid density in regions more distant from the origin, thereby achieving equivalent approximation quality with better efficiency.
Runtime of RA-Value Iteration using $\alpha_{T\epsilon}$.
The runtime of RA-Value Iteration on our running example would be
$O\left(|\mathcal{S}||\mathcal{A}|(T/\alpha_{T\epsilon})^{d}\right)$ $=O\left(|\mathcal{S}||\mathcal{A}|\left(\frac{T^{2}\cdot d^{2}}{\epsilon}\right)^{d}\right).$
It may be helpful to note that, if one takes $|\mathcal{A}|$, the maximum number of actions available from a given state, the discount factor $\gamma$, and the dimension of reward $d$ to be small constants close to 1, then the asymptotic dependence of the runtime on the two fundamental parameters of the size of the state space $|\mathcal{S}|$ and the desired approximation $\epsilon$ is
$O\left(|\mathcal{S}|\left(\frac{1}{\epsilon}\log^{2}(1/\epsilon)\right)^{d}\right)$
for our running example.
Model Learning and and $m_{known}$.
The overall RA-E3 algorithm must learn a model during exploration, unlike the RA-Value Iteration that takes a model as input. $m_{known}$ is the number of times we must visit a state in order to “know” its statistics (especially transition probabilities) sufficiently well. In the worst case bound we give that
$m_{known}=O(((|\mathcal{S}||\mathcal{A}|TG^{T}_{\max})/\epsilon)^{4}|\mathcal{A}|\log(1/\beta))$
but this is extremely pessimistic, as it assumes that every state may transition to every other possible state, and that every action is available in every state, and that there are no independence assumptions between any states, actions, and transition probabilities. In other words, this is the result with no assumptions about the model structure. In practice, for an environment where only a small constant number of actions and transitions are available from a given state, many of the terms disappear.
Appendix B Additional Details on The Analysis of The RA-E3 Algorithm
In this section, we comment on some additional details of the analysis of Theorem 13, the main theorem for the RA-E3 algorithm.
We begin by defining approximation of MOMDPs. Think of $M$ as the true MOMDP, that is, a perfect model of transition probabilities. Think of $M^{\prime}$, on the other hand, as the current best estimate of an MOMDP obtained through exploration. In particular, $M^{\prime}$ in the RA-E3 algorithm will consist of the set of known states.
Definition 14.
(Kearns and Singh 2002) Let $M$ and $M^{\prime}$ be two MOMDPs over the same state space $\mathcal{S}$ with the same deterministic reward function $\mathbf{R}(s,a)$. $M^{\prime}$ is an $\alpha-$approximation of $M$ if for any state $s$ and $s^{\prime}$ and any action $a$,
$Pr_{M}(s^{\prime}\mid s,a)-\alpha\leq Pr_{M^{\prime}}(s^{\prime}\mid s,a)\leq Pr_{M}(s^{\prime}\mid s,a)+\alpha,$
where the subscript denotes the model.
We now extend the Simulation Lemma (Kearns and Singh 2002), which tells us how close the approximation of an MOMDP needs to be in order for the expected welfare, or ESR, of a policy to be close in an estimated model. The argument is similar to (Kearns and Singh 2002), but in our case the policy may not be stationary and we want to bound the deviation in expected welfare rather than just accumulated reward.
Recall that $G^{T}_{max}$ is defined to be the maximum possible welfare achieved on a $T$-step trajectory - $G^{T}_{max}$ is at most $W(\mathbf{T})$ where $\mathbf{T}$ equals $T$ times the identity vector, in our model.
Lemma 15 (Extended Simulation Lemma).
Let $M^{\prime}$ be an $O((\epsilon/|\mathcal{S}||\mathcal{A}|TG^{T}_{\max})^{2})$-approximation of $M$. Then for any policy $\pi$, any state $s$, and horizon time $T$, we have
$V^{\pi}_{M}(s,\tau_{0;0},T)-\epsilon\leq V^{\pi}_{M^{\prime}}(s,\tau_{0;0},T)\leq V^{\pi}_{M}(s,\tau_{0;0},T)+\epsilon.$
Proof.
Fix a policy $\pi$ and a start state $s$. Let $M^{\prime}$ be an $\alpha$-approximation of $M$ (we will later show that $\alpha$ has the same bound as the Lemma statement). Call the transition probability from a state $s$ to a state $s^{\prime}$ under action $a$ to be $\beta$-small in $M$ if $Pr_{M}(s^{\prime}\mid s,a)\leq\beta$. Then the probability that a $T$-trajectory $\tau$ starting from a state $s$ following policy $\pi$ contains at least one $\beta$-small transition is at most $\beta|\mathcal{S}||\mathcal{A}|T$. This is because the total probability of all $\beta$-small transitions in $M$ is at most $\beta|\mathcal{S}||\mathcal{A}|$ (assuming all transition probabilities are $\beta$-small), and there are $T$ timesteps. Note that in our case, the optimal policy may not be necessarily stationary, thus the agent does not necessarily choose the same action (and hence the same transition probability) upon revisiting any state. So we cannot bound the total probability by $\beta|\mathcal{S}|$ like in the original proof.
The total expected welfare of the trajectories of $\pi$ that consist of at least one $\beta$-small transition of $M$ is at most $\beta|\mathcal{S}||\mathcal{A}|TG^{T}_{\max}$. Recall that $M^{\prime}$ be an $\alpha$-approximation of $M$. Then for any $\beta$-small transition in $M$, we have $Pr_{M^{\prime}}(s^{\prime}\mid s,a)\leq\alpha+\beta$. So the total welfare of the trajectories of $\pi$ that consist of at least one $\beta$-small transition of $M$ is at most $(\alpha+\beta)|\mathcal{S}||\mathcal{A}|TG^{T}_{\max}$. We can thus bound the difference between $V^{\pi}_{M}(s,\tau_{0;0},T)$ and $V^{\pi}_{M^{\prime}}(s,\tau_{0;0},T)$ restricted to these trajetories by $(\alpha+2\beta)|\mathcal{S}||\mathcal{A}|TG^{T}_{\max}$. We will later choose $\beta$ and bound this value by $\epsilon/4$ to solve for $\alpha$.
Next, consider trajectories of length $T$ starting at $s$ that do not contain any $\beta$-small transitions, i.e. $Pr_{M}(s^{\prime}\mid s,a)>\beta$ in these trajectories. Choose $\Delta=\alpha/\beta$, we may write
$(1-\Delta)Pr_{M}(s^{\prime}\mid s,a)$
$\qquad\leq Pr_{M^{\prime}}(s^{\prime}\mid s,a)$
$\qquad\leq(1+\Delta)Pr_{M}(s^{\prime}\mid s,a).$
because $M^{\prime}$ is an $\alpha$-approximation of $M$ and $\beta\leq 1$. Thus for any $T$-trajectory $\tau$ that does not cross any $\beta$-small transitions under $\pi$, we have
$(1-\Delta)^{T}Pr_{M}^{\pi}[\tau]\leq Pr_{M^{\prime}}^{\pi}[\tau]\leq(1+\Delta)^{T}Pr_{M}^{\pi}[\tau],$
which follows from the definition of the probability along a trajectory and the fact that $\pi$ is the same policy in all terms. Since we assume reward functions are deterministic in our case, for any particular $T$-trajectory $\tau$, we also have
$W_{M}(\mathbf{R}(\tau))=W_{M^{\prime}}(\mathbf{R}(\tau)).$
Since these hold for any fixed $T$-trajectory that does not traverse any $\beta$-small transitions in $M$ under $\pi$, they also hold when we take expectations over the distributions over such $T$-trajectories in $M$ and $M^{\prime}$ induced by $\pi$. Thus
$(1-\Delta)^{T}V^{\pi}_{M}(s,\tau_{0;0},T)-\frac{\epsilon}{4}$
$\qquad\leq V^{\pi}_{M^{\prime}}(s,\tau_{0;0},T)$
$\qquad\leq(1+\Delta)^{T}V^{\pi}_{M}(s,\tau_{0;0},T)+\frac{\epsilon}{4}.$
where the $\epsilon/4$ terms account for the contributions of the $T$trajectories that traverse at least one $\beta$-small transitions under $\pi$. It remains to show how to choose $\alpha$, $\beta$, and $\Delta$ to obtain the desired approximation. For the upper bound, we solve for
$(1+\Delta)^{T}V^{\pi}_{M}(s,\tau_{0;0},T)\leq V^{\pi}_{M}(s,\tau_{0;0},T)+\frac{\epsilon}{4}$
$\implies(1+\Delta)^{T}\leq 1+\epsilon/(4G^{T}_{\max}).$
By taking $\log$ on both sides and using Taylor expansion, we can upper bound $\Delta$.
$T\Delta/2\leq\epsilon/(4G^{T}_{\max})$
$\qquad\implies\Delta\leq\epsilon/(2TG^{T}_{\max}).$
Choose $\beta=\sqrt{\alpha}$. Then
$\begin{cases}(\alpha+2\beta)|\mathcal{S}||\mathcal{A}|TG^{T}_{\max}\leq 3\sqrt{\alpha}|\mathcal{S}||\mathcal{A}|TG^{T}_{\max}\leq\epsilon/4\\ \Delta=\sqrt{\alpha}\leq\epsilon/(2TG^{T}_{\max})\end{cases}$
Choosing $\alpha=O((\epsilon/|\mathcal{S}||\mathcal{A}|TG^{T}_{\max})^{2})$ solves the system. The lower bound can be handled similarly, which completes the proof of the lemma. ∎
We now define a “known state.” This is a state that has been visited enough times, and its actions have been trialed sufficiently many times, that we have accurate estimates of the transition probabilities from this state.
Definition 16.
(Kearns and Singh 2002) Let $M$ be an MOMDP. A state $s$ of $M$ is considered known if it has be visited a number of times equal to
$m_{known}=O((|\mathcal{S}||\mathcal{A}|TG^{T}_{\max}/\epsilon)^{4}|\mathcal{A}|\log(1/\delta)).$
By applying Chernoff bounds, we can show that if a state has been visited $m_{known}$ times then its empircal estimation of the transition probabilities satisfies the accuracy required by the Lemma 15.
Lemma 17.
(Kearns and Singh 2002) Let $M$ be an MOMDP. Let $s$ be a state of $M$ that has been visited at least $m$ times, with each action having been executed at least $\lfloor m/|\mathcal{A}|\rfloor$ times. Let $\hat{Pr}(s^{\prime}\mid s,a)$ denote the empirical transition probability estimates obtained from the $m$ visits to $s$. Then if
$m=O((|\mathcal{S}||\mathcal{A}|TG^{T}_{\max}/\epsilon)^{4}|\mathcal{A}|\log(1/\delta)),$
then with probability $1-\delta$, we have
$|\hat{Pr}(s^{\prime}\mid s,a)-Pr(s^{\prime}\mid s,a)|=O((\epsilon/|\mathcal{S}||\mathcal{A}|TG^{T}_{\max})^{2})$
for all $s^{\prime}\in\mathcal{S}$.
Proof.
The sampling version of Chernoff bounds states that if the number of independent, uniformly random samples $n$ that we use to estimate the fraction of a population with certain property $p$ satisfies
$n=O\left(\frac{1}{\alpha^{2}}\log\left(\frac{1}{\delta}\right)\right),$
the our estimate $\bar{X}$ satisfies
$\bar{X}\in[p-\alpha,p+\alpha]~{}\text{with probability $1-\delta$}.$
By the Extended Simulation Lemma, it suffices to choose $\alpha=O((\epsilon/|\mathcal{S}||\mathcal{A}|TG^{T}_{\max})^{2})$.
Note that we need to insert an extra factor of $|\mathcal{A}|$ compared to the original analysis since we treat the size of the action space as a variable instead of a constant, and a state is categorized as “known” only if the estimates of transition probability of all actions are close enough. ∎
We have specified the degree of approximation required for sufficient simulation accuracy. It remains to directly apply the Explore or Exploit Lemma from (Kearns and Singh 2002) to conclude the analysis.
Lemma 18 (Explore or Exploit Lemma (Kearns and Singh 2002)).
Let $M$ be any MOMDP, let $S$ be any subset of the states of $M$, and let $M_{S}$ be the MOMDP on $M$. For any $s\in\mathcal{S}$, any $T$, and any $01>\alpha>$ 0, either there exists a policy $\pi$ in $M_{S}$ such that $V^{\pi}_{M_{S}}(s,\tau_{0;0},T)\geq V^{*}_{M}(s,\mathbf{0},T)-\alpha$, or there exists a policy $\pi$ in $M_{S}$ such that the probability that a $T$-trajectory following $\pi$ will lead to the exit state exceeds $\alpha/G^{T}_{\max}$.
This lemma guarantees that either the $T$step return of the optimal exploitation policy in the simulated model is very close to the optimal achievable in $M$, or the agent choosing the exploration policy can reach a previously unknown state with significant probability.
Note that unlike the original E3 algorithm, which uses standard value iteration to compute the exactly optimal policy (optimizing a linear function of a scalar reward) on the sub-model $\hat{M}_{S}$, we use our RA-Value Iteration algorithm to find an approximately optimal policy. Therefore, we need to allocate $\epsilon/2$ error to both the simulation stage and the exploitation stage, which gives a total of $\epsilon$ error for the entire algorithm.
It remains to handle the failure parameter $\beta$ in the statement of the main theorem, which can be done similarly to (Kearns and Singh 2002). There are two sources of failure for the algorithm:
  • The algorithm’s estimation of the next state distribution for some action at a known state is inaccurate, resulting $\hat{M}_{S}$ being an inaccurate model of $M_{S}$.
  • Despite doing attempted explorations repeatedly, the algorithm fails to turn a previously unknown state into a known state because of an insufficient number of balanced wandering.
It suffices to allocate $\beta/2$ probabilities to each source of failure. The first source of failure are bounded by Lemma 17. By choosing $\beta^{\prime}=\beta/(2|\mathcal{S}|)$, we ensure that the probability that the first source of failure happens to each of the known states in $M_{S}$ is sufficiently small such that the total failure probability is bounded by $\beta/2$ for $M_{S}$.
For the second source of failure, by the Explore or Exploit Lemma, each attempted exploration results in at least one step of balanced wandering with probability at least $\epsilon/(2G^{T}_{\max})$, i.e. when this leads the agent to an unknown state. The agent does at most $|\mathcal{S}|m_{known}$ steps of balanced wandering, since this makes every state known. By Chernoff bound, the probability that the agent does fewer than $|\mathcal{S}|m_{known}$ steps of balanced wandering (attempted explorations that actually leads to an unknown state) will be smaller than $\beta/2$ if the number of attempted explorations is
$O((G^{T}_{\max}/\epsilon)|\mathcal{S}|\log(|\mathcal{S}|/\beta)m_{known}),$
where $m_{known}=O(((|\mathcal{S}||\mathcal{A}|TG^{T}_{\max})/\epsilon)^{4}|\mathcal{A}|\log(|\mathcal{S}|/\beta))$ (recall we choose $\beta^{\prime}=\beta/(2|\mathcal{S}|)$ for the first source of failure).
Thus, the total computation time is bounded by $O\left(|\mathcal{S}||\mathcal{A}|(T/\alpha_{T\epsilon/2})^{d}\right)$ (the time required for RA-Value Iteration during off-line computations with $\epsilon/2$ precision by Lemma 11) times $T$ times the maximum number of attempted explorations, giving
$O((|\mathcal{S}|^{2}|\mathcal{A}|TG^{T}_{\max}/\epsilon)(T/\alpha_{T\epsilon/2})^{d}\log(|\mathcal{S}|/\beta)m_{known}).$
This concludes the remaining details of the analysis of the RA-E3 algorithm.