### v3

# Multi-objective Reinforcement Learning with Nonlinear Preferences: Provable Approximation for Maximizing Expected Scalarized Return

Abstract
We study multi-objective reinforcement learning with nonlinear preferences over trajectories. That is, we maximize the expected value of a nonlinear function over accumulated rewards (expected scalarized return or ESR) in a multi-objective Markov Decision Process (MOMDP). We derive an extended form of Bellman optimality for nonlinear optimization that explicitly considers time and current accumulated reward. Using this formulation, we describe an approximation algorithm for computing an approximately optimal non-stationary policy in pseudopolynomial time for smooth scalarization functions with a constant number of rewards. We prove the approximation analytically and demonstrate the algorithm experimentally, showing that there can be a substantial gap between the optimal policy computed by our algorithm and alternative baselines.
1 Introduction
Markov Decision Processes (MDPs) model goal-driven interaction with a stochastic environment, typically aiming to maximize a scalar-valued reward per time-step through a learned policy. Equivalently, this formulation asks the agent to maximize the expected value of a linear function of total reward. This problem can be solved with provable approximation guarantees in polynomial time with respect to the size of the MDP [16, 9, 5, 6, 2].
We extend this framework to optimize a nonlinear function of vector-valued rewards in multi-objective MDPs (MOMDPs), aiming to maximize expected welfare $\mathbb{E}[W(\mathbf{r})]$ where $\mathbf{r}$ is a total reward vector for $d$ objectives. We note that this function is also called the utility or the scalarization function within the multi-objective optimization literature [14]. Unlike deep neural networks used for large state spaces which uses nonlinear function approximators, our nonlinearity lies entirely in the objective function.
Motivation.
Nonlinear welfare functions capture richer preferences for agents, such as fairness or risk attitudes. For example, the Nash Social Welfare function $\left(W_{\text{Nash}}(\mathbf{r})=\left(\prod_{i}r_{i}\right)^{1/d}\right)$ reflects a desire for fairness or balance across objectives and diminishing marginal returns [11]. Even single-objective decision theory uses nonlinear utility functions to model risk preferences, like risk aversion with the Von Neumann-Morgenstern utility function [24].
Consider a minimal example where an autonomous taxi robot, Robbie, serving rides in neighborhoods $A$ and $B$, as diagrammed in Figure 1. Each ride yields a reward in its respective dimension.
$A$$B$$(1,0)$$(0,1)$$(0,0)$
Figure 1: Taxi Optimization Example
Suppose Robbie starts in neighborhood $A$ and has $t=3$ time intervals remaining before recharging. With no discounting, the Pareto Frontier of maximal (that is, undominated) policies can achieve cumulative reward vectors of $(3,0)$, $(1,1)$, or $(0,2)$. If we want Robbie to prefer the second more balanced or “fair” option of serving one ride in each of the two neighborhoods then Robbie must have nonlinear preferences. However, the second option would maximize the Nash Social Welfare of cumulative reward, for example.
Nonlinear preferences complicate policy computation: Bellman optimality fails, and stationary policies may be suboptimal. Intuitively, a learning agent with fairness oriented preferences to balance objectives should behave differently, even in the same state, depending on which dimension of reward is “worse off.” In the Figure 1 example, one policy to achieve the balanced objective $(1,1)$ is to complete a ride in $A$, then travel from $A$ to $B$, and finally to complete a ride in $B$ – note that this is not stationary with respect to the environment states.
Contributions.
In this work, we ask whether it is possible to describe an approximation algorithm (that is, with provable guarantees to approximate the welfare optimal policy) for MOMDPs with nonlinear preferences that has polynomial dependence on the number of states and actions, as is the case for linear preferences or scalar MDPs. To the best of our knowledge, ours is the first work to give provable guarantees for this problem, compared to other work that focuses exclusively on empirical evaluation of various neural network architectures.
We show this is possible for smooth preferences and a constant number of dimensions of reward. To accomplish this, we (i) derive an extended form of Bellman optimality (which may be of independent interest) that characterizes optimal policies for nonlinear preferences over multiple objectives, (ii) describe an algorithm for computing approximately optimal non-stationary policies, (iii) prove the worst-case approximation properties of our algorithm, and (iv) demonstrate empirically that our algorithm can be used in large state spaces to find policies that significantly outperform other baselines.
2 Related Work
Most algorithms for reinforcement learning focus on a single scalar-valued objective and maximizing total expected reward [21]. Classic results on provable approximation and runtime guarantees for reinforcement learning include the E3 algorithm [16]. This result showed that general MDPs could be solved to near-optimality efficiently, meaning in time bounded by a polynomial in the size of the MDP (number of states and actions) and the horizon time. Subsequent results refined the achievable bounds [9, 5, 6]. We extend these results to the multi-objective case with nonlinear preferences.
Multi-objective reinforcement learning optimizes multiple objectives at once. So-called single-policy methods use a scalarization function to reduce the problem to scalar optimization for a single policy, and we follow this line of research. The simplest form is linear scalarization, applying a weighted sum on the Q vector [18, 17, 1, 25, 4].
A more general problem is to optimize the expected value of a potentially nonlinear function of total reward, which may be vector-valued for a multi-objective optimization context. We refer to such a function as a welfare function [7, 13, 20]; it is also commonly referred to as a utility or scalarization function [14, 3]. Other recent work has studied nonlinear objectives, but ours is the first to our knowledge to provide an approximation algorithm with provable guarantees based on a characterization of recursive optimality in this setting. Several others study algorithms with desirable convergenece properties and empirical performance based on conditioning function approximators on accumulated reward but with no approximation guarantees, all of which we see as complementary efforts to our work [20, 13, 10, 19].
[3] describe another model-based algorithm that can compute approximately optimal policies for a general class of monotone and Lipshitz welfare functions, but rather than maximizing the expected welfare, they maximize the welfare of expected rewards (note the two are not equal for nonlinear welfare functions). Other works have formulated fairness in different ways or settings. For example, [15] defines an analogue of envy freeness and [12] studies a per-time-step fairness guarantee. [7] studies welfare maximization in multi-armed bandit problems.
3 Preliminaries
A Multi-objective Markov Decision Process (MOMDP) consists of a finite set $\mathcal{S}$ of states, a starting state $s_{1}\in\mathcal{S}$, [^1] In general we may have a distribution over starting states; we assume a single starting state for ease of exposition. a finite set $\mathcal{A}$ of actions, and probabilities $Pr(s^{\prime}|s,a)\in[0,1]$ that determine the probability of transitioning to state $s^{\prime}$ from state $s$ after taking action $a$. Probabilities are normalized so that $\sum_{s^{\prime}}Pr(s^{\prime}|s,a)=1$.
We also have a vector-valued reward function $\mathbf{R}(s,a):\mathcal{S}\times\mathcal{A}\rightarrow[0,1]^{d}$. Each of the $d$ dimensions of the reward vector corresponds to one of the multiple objectives that are to be maximized. At each time step $t$, the agent observes state $s_{t}\in\mathcal{S}$, takes action $a_{t}\in\mathcal{A}(s_{t})$, and receives a reward vector $\mathbf{R}(s_{t},a_{t})\in\mathbb{R}^{d}$. The environment, in turn, transitions into $s_{t+1}$ with probability $Pr(s_{t+1}|a_{t},s_{t})$.
To make the optimization objective well-posed in MOMDPs with vector-valued rewards, we must specify a scalarization function [14] which we denote as $W:\mathbb{R}^{d}\to\mathbb{R}$. For fair multi-objective reinforcement learning, we think of each of the $d$ dimensions of the reward vector as corresponding to distinct users. The scalarization function can thus be thought of as a welfare function over the users, and the learning agent is a welfare maximizer. Even when $d=1$, a nonlinear function $W$ can be a Von Neumann-Morgenstern utility function [24] that expresses the risk-attitudes of the learning agent, with strictly concave functions expressing risk aversion.
Assumptions.
Here we clarify some preliminary assumptions we make about the reward function $\mathbf{R}(s,a):\mathcal{S}\times\mathcal{A}\rightarrow[0,1]^{d}$ and the welfare function $W:\mathbb{R}^{d}\to\mathbb{R}$. The restriction to $[0,1]$ is not too restrictive and can be seen simply as a normalization; the substantial assumption is that the rewards are simply finite and bounded. Note from the writing we assume the reward function is deterministic: While it suffices for linear optimization to simply learn the mean of random reward distributions, this does not hold when optimizing the expected value of a nonlinear function. Nevertheless, the environment itself may still be stochastic, as the state transitions may still be random.
We also assume that $W$ is smooth. For convenience of analysis, we will assume $W$ is uniformly continuous on the L1-norm (other parameterizations such as the stronger Lipshitz continuity or using the L2-norm are also possible). For all $\epsilon>0$, there exists $\delta_{\epsilon}$ such that for all $\mathbf{x},\mathbf{y}\geq 0$,
$|W(\mathbf{x})-W(\mathbf{y})|<\epsilon\quad\text{if}\quad\|\mathbf{x}-\mathbf{y}\|_{1}<\delta_{\epsilon}.$
This smoothness assumption is necessary for worst-case analysis, but not to run the algorithm.
4 Modeling Optimality
In this section, we expand the classic model of reinforcement learning to optimize the expected value of a nonlinear function of (possibly) multiple dimensions of reward. We begin with the notion of a trajectory of state-action pairs.
Definition 1.
Let $M$ be an MOMDP. A length $T$ trajectory in $M$ is a tuple $\tau$ of $T$ state-action pairs
$(s_{1},a_{1}),(s_{2},a_{2}),\dots,(s_{T},a_{T}).$
For $1\leq k\leq k^{\prime}\leq T$, let $\tau_{k:k^{\prime}}$ be the sub-trajectory consisting of pairs $(s_{k},a_{k}),\dots,(s_{k^{\prime}},a_{k^{\prime}})$. Let $\tau_{0;0}$ denote the empty trajectory.
For a discount factor $\gamma$, we calculate the total discounted reward of a trajectory. Note that this is a vector in general.
Definition 2.
For length $T$ trajectory $\tau$ and discount factor $10\leq\gamma<$ 1, the total discounted reward along $\tau$ is the vector
$\mathbf{R}(\tau,\gamma)=\sum_{t=1}^{T}\gamma^{t-1}\mathbf{R}(s_{t},a_{t}).$
For ease of exposition we will frequently leave $\gamma$ implicit from context and simply write $\mathbf{R}(\tau)$.
A policy is a function $\pi(a\mid\tau,s)\in[0,1]$ mapping past trajectories and current states to probability distributions over actions, that is, $\sum_{a}\pi(a\mid\tau,s)=1$ for all $\tau$ and $s$. A stationary policy is the special case of a policy that depends only on the current state: $\pi(a\mid s)$.
Definition 3.
The probability that a $T$-trajectory $\tau$ is traversed in an MOMDP $M$ upon starting in state $s_{1}$ and executing policy $\pi$ is
$Pr^{\pi}[\tau]=\pi(a_{1}\mid\tau_{0:0},s_{1})\times\prod_{t=2}^{T}\pi\left(a_{t}~{}\bigg{|}~{}\tau_{1:t-1},s_{t}\right)Pr(s_{t}|s_{t-1},a_{t-1}).$
Problem Formulation.
Given a policy, a finite time-horizon $T$, and a starting state $s_{1}$ we can calculate the expected welfare of total discounted reward along a trajectory as follows. Our goal is to maximize this quantity. That is, we want to compute a policy that maximizes the expected $T$-step discounted welfare.
Definition 4.
For a policy $\pi$ and a start state $s$, the expected $T$-step discounted welfare is
$\mathbb{E}_{\tau\sim\pi}\bigl{[}W\bigl{(}\mathbf{R}(\tau)\bigr{)}\bigr{]}=\sum_{\tau}Pr^{\pi}[\tau]W(\mathbf{R}(\tau))$
where the expectation is taken over all length $T$ trajectories beginning at $s$.
Note that this objective is not equal to $W\left(\mathbb{E}_{\tau\sim\pi}\bigl{[}\mathbf{R}(\tau)\bigr{]}\right)$, which others have studied [3, 20], for a nonlinear $W$. The former (our objective) is also known as expected scalarized return (ESR) whereas the latter is also known as scalarized expected return (SER) [14]. While SER makes sense for a repeated decision-making problem, it does not optimize for expected welfare for any particular trajectory. For concave $W$, ESR $\leq$ SER by Jensen’s inequality. However, an algorithm for approximating SER does not provide any guarantee for approximating ESR. For example, a policy can be optimal on SER but achieve 0 ESR if it achieves high reward on one or the other of two objectives but never both in the same episode.
Form of Optimal Policy and Value Functions.
The optimal policy for this finite time horizon setting is a function also of the number of time steps remaining. We write such a policy as $\pi(a\mid s,\tau,t)$ where $\tau$ is a trajectory (the history), $s$ is the current state, and $t$ is the number of time steps remaining, $i.e.$ $t=T-|\tau|$.
We can similarly write the extended value function of a policy $\pi$. We write $\tau$ as the history or trajectory prior to some current state $s$ and $\tau^{\prime}$ as the future, the remaining $t$ steps determined by the policy $\pi$ and the environmental transitions.
Definition 5.
The value of a policy $\pi$ beginning at state $s$ after history $\tau$ and with $t$ more actions is
$V^{\pi}(s,\tau,t)=\mathbb{E}_{\tau^{\prime}\sim\pi}\left[W(\mathbf{R}(\tau)+\gamma^{T-t}\mathbf{R}(\tau^{\prime}))\right]$
where the expectation is taken over all length $t$ trajectories $\tau^{\prime}$ beginning at $s$. The optimal value function is
$V^{*}(s,\tau,t)=\max_{\pi}V^{\pi}(s,\tau,t)$
and the optimal policy is
$\pi^{*}\in\operatorname*{arg\,max}_{\pi}V^{\pi}(s,\tau,t).$
Note that because $W$ is nonlinear, the value function $V^{\pi}$ cannot be decomposed in the same way as in the traditional Bellman equations. Before proceeding we want to provide some intuition for this point. The same reasoning helps to explain why stationary policies are not generally optimal.
Suppose we are optimizing the product of the reward between two objectives (i.e., the welfare function is the product or geometric mean), and at some state $s$ with some prior history $\tau$ we can choose between two policies $\pi_{1}$ or $\pi_{2}$. Suppose that the future discounted reward vector under $\pi_{1}$ is $(1,1)$, whereas it is $(10,0)$ under $\pi_{2}$. So $\pi_{1}$ has greater expected future welfare and traditional Bellman optimality would suggest we should choose $\pi_{1}$. However, if $\mathbf{R}(\tau)=(0,10)$, we would actually be better off in terms of total welfare choosing $\pi_{2}$. In other words, both past and future reward are relevant when optimizing for the expected value of a nonlinear welfare function.
We develop an extended form of Bellman optimality capturing this intuition by showing that the optimal value function can be written as a function of current state $s$, accumulated discounted reward $\mathbf{R}(\tau)$, and number of timesteps remaining in the task $t$. The proof is deferred to the appendix.
Lemma 1.
Let $\mathcal{V}(s,\mathbf{R}(\tau),0)=W(\mathbf{R}(\tau))$ for all states $s$ and trajectories $\tau$. For every state $s$, history $\tau$, and $t>0$ time steps remaining, let
$\mathcal{V}(s,\mathbf{R}(\tau),t)=\max_{a}\mathbb{E}_{s^{\prime}}\left[\mathcal{V}(s^{\prime},\mathbf{R}(\tau)+\gamma^{T-t}\mathbf{R}(s,a),t-1)\right].$
Then $V^{*}(s,\tau,t)=\mathcal{V}(s,\mathbf{R}(\tau),t)$.
By Lemma 1, we can parameterize the optimal value function by the current state $s$, accumulated reward $\mathbf{R_{acc}}\in\mathbb{R}^{d}$, and number of timesteps remaining $t$. We will use this formulation of $V^{*}$ in the remainder of the paper.
Definition 6 (Recursive Formulation of $V^{*}$).
Let $\mathbf{R_{acc}}=\mathbf{R}(\tau)$ be the vector of accumulated reward along a history prior to some state $s$. Let $V^{*}(s,\mathbf{R_{acc}},0)=W(\mathbf{R_{acc}})$ for all $s$ and $\mathbf{R_{acc}}$. For $t>0$,
$V^{*}(s,\mathbf{R}_{acc},t)=\max_{a}\sum_{s^{\prime}}Pr(s^{\prime}|s,a)\times V^{*}(s^{\prime},\mathbf{R}_{acc}+\gamma^{T-t}\mathbf{R}(s,a),t-1).$
Horizon Time.
Note that an approximation algorithm for this discounted finite time horizon problem can also be used as an approximation algorithm for the discounted infinite time horizon problem. Informally, discounting by $\gamma$ with bounded maximum reward implies that the first $T\approx 1/(1-\gamma)$ steps dominate overall returns. We defer the precise formulation of the lower bound on the horizon time and its proof in the appendix.
5 Computing Optimal Policies
Our overall algorithm is Reward-Aware Explore or Exploit (or RAEE for short) to compute an approximately optimal policy, inspired by the classical E3 algorithm [16]. At a high level, the algorithm explores to learn a model of the environment, periodically pausing to recompute an approximately optimal policy on subset of the environment that has been thoroughly explored. We call this optimization subroutine Reward-Aware Value Iteration (or RAVI for short). In both cases, reward-aware refers to the fact that the algorithms compute non-stationary policies in the sense that optimal behavior depends on currently accumulated vector-valued reward, in addition to the current state.
Of the two, the model-based optimization subroutine RAVI is the more significant. The integrated model-learning algorithm RAEE largely follows from prior work, given access to the RAVI subroutine. For this reason, we focus in this section on RAVI, and defer a more complete discussion and analysis of RAEE to Section 7 and the appendix.
A naive algorithm for computing a nonstationary policy would need to consider all possible prior trajectories for each decision point, leading to a runtime complexity containing the term $|\mathcal{S}|^{T}$, exponential in the size of the state space and time horizon. Instead, for a smooth welfare function on a constant number of objectives, our algorithm will avoid any exponential dependence on $|\mathcal{S}|$.
The algorithm is derived from the recursive definition of the optimal multi-objective value function $V^{*}$ in Definition 6, justified by Lemma 1, parameterized by the accumulated discounted reward vector $\mathbf{R_{acc}}$ instead of the prior history. Note that even if rewards were integer, $\mathbf{R_{acc}}$ might not be due to discounting. We must therefore introduce a discretization that maps accumulated reward vectors to points on a lattice, parameterized by some $\alpha\in(0,1)$ where a smaller $\alpha$ leads to a finer discretization but increases the runtime.
Definition 7.
For a given discretization precision parameter $\alpha\in(0,1)$, define $f_{\alpha}:\mathbb{R}^{d}\to(\alpha\mathbb{Z})^{d}$ by
$f_{\alpha}(\mathbf{R})=\left(\left\lfloor\frac{R_{1}}{\alpha}\right\rfloor\cdot\alpha,\left\lfloor\frac{R_{2}}{\alpha}\right\rfloor\cdot\alpha,\cdots,\left\lfloor\frac{R_{d}}{\alpha}\right\rfloor\cdot\alpha\right).$
In other words, $f$ maps any $d$-dimension vector to its nearest $\alpha$ multiples. We are now prepared to describe the algorithm, which at a high level, uses computes the dynamic program of approximately welfare-optimal value functions conditioned on discretized accumulated reward vectors.
Algorithm 1 Reward-Aware Value Iteration (RAVI)
1:Parameters: Discretization precision $\alpha\in(0,1)$, Discount factor $\gamma\in[0,1)$, Reward dimension $d$, finite time horizon $T$, welfare function $W$, discretization function $f_{\alpha}$.
2:Require: Normalize $\mathbf{R}(s,a)\in[0,1]^{d}$ for all $s\in\mathcal{S},a\in\mathcal{A}$.
3:Initialize: $V(s,\mathbf{R_{acc}},0)=W(\mathbf{R_{acc}})$ for all $\mathbf{R_{acc}}\in\{0,\mathbf{\alpha},2\mathbf{\alpha},\dots,\lceil T/\alpha\rceil\cdot\alpha\}^{d}$, $s\in\mathcal{S}$.
4:for t = 1 to $T$ do
5:     for $\mathbf{R_{acc}}\in\{0,\mathbf{\alpha},2\mathbf{\alpha},\dots,\lceil(T-t)/\alpha\rceil\cdot\alpha\}^{d}$ do
6:         for all $s\in\mathcal{S}$ do
7:
$V(s,\mathbf{R_{acc}},t)\leftarrow\max_{a}\sum_{s^{\prime}}Pr(s^{\prime}|s,a)\times V(s^{\prime},f_{\alpha}(\mathbf{R}_{acc}+\gamma^{T-t}\mathbf{R}(s,a)),t-1)$
8:
$\pi(s,\mathbf{R_{acc}},t)\leftarrow\operatorname*{arg\,max}_{a}\sum_{s^{\prime}}Pr(s^{\prime}|s,a)\times V(s^{\prime},f_{\alpha}(\mathbf{R}_{acc}+\gamma^{T-t}\mathbf{R}(s,a)),t-1)$
The asymptotic runtime complexity is $O\left(|\mathcal{S}|^{2}|\mathcal{A}|(T/\alpha)^{d}\right)$. However, observe that the resulting algorithm is extremely parallelizable: Given solutions to subproblems at $t-1$, all subproblems for $t$ can in principle be computed in parallel. .
The parallel computing implementation of the RAVI algorithm leverages the computational power of GPUs to handle the extensive calculations involved in multi-objective value iteration. Each thread computes the value updates and policy decisions for a specific state and accumulated reward combination, which allows for massive parallelism.
In practice, we observed an empirical speedup of approximately 560 times on a single NVIDIA A100 compared to the CPU implementation on an AMD Ryzen 9 7950x 16-core processor for our experimental settings. This drastic improvement in runtime efficiency makes it feasible to run RAVI on larger and more complex environments, where the computational demands would otherwise be prohibitive.
It remains to see how small $\alpha$ needs to be, which will dictate the final runtime complexity. We first analyze the correctness of the algorithm and then return to the setting of $\alpha$ for a given smoothness of $W$ to conclude the runtime analysis.
We begin analyzing the approximation of RAVI with an important structural property: the optimal value function is smooth if the welfare function is smooth. The proof is included in the appendix.
Lemma 2 (Uniform continuity of multi-objective value function).
Let the welfare function $W:\mathbb{R}^{d}\to\mathbb{R}$ be uniformly continuous. Fix $s\in\mathcal{S}$ and $t\in\{0,1,\dots,T\}$, then for all $\epsilon>0$, there exists $\delta_{\epsilon}>0$ such that
$\bigg{|}V^{*}(s,\mathbf{R_{1}},t)-V^{*}(s,\mathbf{R_{2}},t)\bigg{|}<\epsilon\quad\text{if}\quad\|\mathbf{R_{1}}-\mathbf{R_{2}}\|<\delta_{\epsilon}.$
We now present the approximation guarantee of RAVI, that it achieves an additive error that scales with the smoothness of $W$ and the number of time steps remaining. The proof of this lemma is deferred to the appendix.
Lemma 3 (Approximation Error of RAVI).
For uniformly continuous welfare function $W$, for all $\epsilon>0$, there exists $\alpha_{\epsilon}$ such that
$V(s,\mathbf{R_{acc}},t)\geq V^{*}(s,\mathbf{R_{acc}},t)-t\epsilon$
$\forall s\in\mathcal{S},\mathbf{R_{acc}}\in\mathbb{R}^{d},t\in\{0,1,\dots,T\}$, where $V(s,\mathbf{R_{acc}},t)$ is computed by Algorithm 1.
While we can use any setting of $\alpha$ empirically, this shows that for an approximation guarantee we should set the discretization parameter to $\alpha=\delta_{T\epsilon}/d$ where $\delta_{T\epsilon}$ from the proof of Lemma 3 is sufficient to drive its bound to $\epsilon$, that is, it should be $\alpha=\frac{\delta_{\epsilon}}{T\cdot d}.$
We thus arrive at the ultimate statement of the approximation and runtime of RAVI. The proof follows directly from Lemma 3 and the setting of $\alpha$.
Theorem 1 (Optimality Guarantee of RAVI).
For a given $\epsilon$ and welfare function $W$ that is $\delta_{\epsilon}$ uniformly continuous, RAVI with $\alpha=\frac{\delta_{\epsilon}}{T\cdot d}$ computes a policy $\hat{\pi}$, such that $V^{\hat{\pi}}(s,0,T)\geq V^{*}(s,0,T)-\epsilon$ in $O\left(|\mathcal{S}|^{2}|\mathcal{A}|(d\cdot T^{2}/\delta_{\epsilon})^{d}\right)$ time.
A concrete example of a particular welfare function, the setting of relevant parameters, and a derivation of a simplified runtime statement may help to clarify Theorem 1. Consider the smoothed proportional fairness objective: $W_{\text{SPF}}(\mathbf{x})=\sum_{i=1}^{d}\ln(x_{i}+1)$, a smoothed log-transform of the Nash Social Welfare (or geometric mean) with better numerical stability.
Uniform continuity is a weaker modeling assumption than $L$-Lipshitz continuity for constant $L$. Note that for an $L$-Lipshitz function, the correct value of $\delta_{\epsilon}$ is just $\epsilon/L$, where $\epsilon$ is the desired approximation factor of the algorithm. By taking the gradient, we can see that the $W_{SPF}(\mathbf{x})$ is $d$-Lipshitz for nonnegative $\mathbf{x}$, where $d$ is the dimension of vector reward, so $\delta_{\epsilon}=\epsilon/d$. In practice, as we will see later, this is actually pessimistic: The objective is significantly smoother farther away from the origin, and a discretization that grows increasingly coarse farther from the origin can still efficiently approximate the objective. Nevertheless, Theorem 1 on this example would give us a runtime of
$O\left(|\mathcal{S}|^{2}|\mathcal{A}|(T/\alpha_{T\epsilon})^{d}\right)=O\left(|\mathcal{S}||\mathcal{A}|\left(\frac{T^{2}\cdot d^{2}}{\epsilon}\right)^{d}\right).$
To further simplify, if one takes the number of actions $|\mathcal{A}|$, discount factor $\gamma$, and dimension $d$ to be small constants, then the asymptotic dependence of the runtime (in our running example) is
$O\left(|\mathcal{S}|^{2}\left(\frac{1}{\epsilon}\log^{2}(1/\epsilon)\right)^{d}\right).$
By contrast, a brute force algorithm that conditions on all previous trajectories, even with these assumptions must scale no better than $O(|\mathcal{S}|^{T})$.
6 Experiments and Empirical Observations
We test RAVI on two distinct interpretations of multi-objective reinforcement learning: A. the fairness interpretation where the agent tries to maximize rewards across all dimensions. B. the objective interpretation, where the agent tries to maximize one while minimizing the other. In both scenarios, we show that RAVI can discover more optimal policies than other baselines for nonlinear multi-objective optimization. This holds even for coarser settings of $\alpha$ in the algorithm than would be necessary for strong theoretical worst-case approximation guarantees.
6.1 Taxi
We use the taxi multi-objective simulation environment considered in [13] for testing nonlinear ESR maximization. In this grid-world environment (Figure 3), the agent is a taxi driver whose goal is to deliver passengers. There are multiple queues of passengers, each having a given pickup and drop-off location. Each queue is a different objective or dimension of the reward. The taxi can only take one passenger at a time and wants to balance trips for passengers in the different queues.
We adopt the Nash social welfare function $W_{\text{Nash}}(\mathbf{r})=\left(\prod_{i}r_{i}\right)^{1/d}$ and egalitarian welfare $W_{\text{Egalitarian}}(\mathbf{r})=\min\{r_{i}\}_{i}$ for fairness considerations. To ensure numerical stability, Nash Welfare is optimized with its smoothed log-transform $W_{\text{SPF}}(\mathbf{r},\lambda)=\sum_{i=1}^{d}\ln(r_{i}+\lambda)$ for $\lambda=10^{-7}$. During evaluations, we still use $W_{\text{Nash}}$.
We consider two intuitive baseline approaches for optimization in a resource allocation task like this one. A. Linearly Scalarized Policy [23]: A relatively straightforward technique for multi-objective RL optimization is to apply the linear combination on the Q-values for each objective. Given weights $\mathbf{w}\in\mathbb{R}^{d}$, $\sum_{i}^{d}w_{i}=1$, the scalarized objective is $SQ(s,a)=\mathbf{w}^{\top}\mathbf{Q}(s,a)$, where $Q(s,a)_{i}$ is the Q-value for the $i^{\text{th}}$ objective. B. Mixture Policy [22]: this baseline works by combining multiple Pareto optimal base policies into a single policy. The mixing is achieved by adopting a base policy for $I$ steps and then switching to the next.
All the hyperparameters, including the interval for mixture policy and the weights for linear scalarization, are tuned via grid search.
We run RAVI along with the baselines on 20 random initializations of a $10\times 10$ taxi environment with three passenger spawn locations and destinations ($d=3$), $T=100$ without reward discount ($\gamma=1$), and compare their scalarized total rewards at convergence. As shown in 1, we observe that RAVI is able to compute better performing policies across both nonlinear objectives compared to both baselines.
6.2 Scavenger Hunt
Inspired by the Resource Gathering environment [8], the scavenger hunt environment is a grid-world simulation where the agent must collect resources while avoiding enemies scattered across the grid. The state representation includes the agent’s position and the status of the resources (collected or not). The reward function is vector-valued, where the first component indicates the number of resources collected, and the second component indicates the damage taken by encountering enemies.
Figure 2: Illustration of the axi environment. The taxi is represented by the yellow circle and passenger spawn and delivery locations are represented by colored cells. Note that the scale of the actual experiments are much larger.
Figure 3: Illustration of the scavenger hunt environment. The agent is represented by the blue cell, the green cells are resources, and the red cells are enemies.
Nash Welfare Egalitarian Welfare RAVI 5.06$\pm$0.13 4.35$\pm$0.49 Mixture Policy 3.99$\pm$0.43 2.86$\pm$0.33 Linearly Scalarized Policy 0.02$\pm$0.26 0.03$\pm$0.14
Table 1: Comparison with baselines in terms of ESR for the taxi environment.
Cobb-Douglas RD Threshold RAVI 1.63$\pm$0.20 3.12$\pm$0.48 Welfare Q-learning 1.044$\pm$0.20 1.33$\pm$0.54 Linearly Scalarized Policy 1.02$\pm$0.32 1.68$\pm$0.50
Table 2: Comparison with baselines in terms of ESR for the scavenger hunt environment.
Given the differing nature of the task from the taxi environment, where the objective is to maximize one reward component while minimizing the other, we adopt two specific scalarization functions for the scavenger hunt environment: A. Resource-Damage (RD) Threshold Scalarization: $RD_{\text{threshold}}(R,D)=R-\max(0,(D-\text{threshold})^{3})$, for the number of resources collected $R$ and the damage taken $D$. The threshold parameter, i.e. the budget, determines the point at which the penalty on the damage taken starts to apply. The RD threshold scalarization encodes the concept of a budget into the optimization process. It is a constrained optimization approach with a soft penalty on the damage taken. B. Cobb-Douglas Scalarization: $CD(R,D)=R^{\alpha}\left(\frac{1}{D+1}\right)^{\beta}$. The Cobb-Douglas scalarization function is inspired by economic theory. It balances the trade-off between the number of resources collected and the damage taken. The parameters $\alpha$ and $\beta$ represent the weights assigned to each component, with $\alpha+\beta=1$. Note that when $\alpha$ and $\beta$ are set to 0.5, this scalarization function reduces $W_{\text{Nash}}(R,1/(D+1))$.
Due to the differing nature of the scavenger hunt task, a mixture policy does poorly — sometimes it ignores the resources and just tries not to hit an enemy, other times it ignores the enemies and just tries to get resources. We therefore compare the performance of RAVI against a linearly scalarized policy and Welfare Q-learning [13]: by using nonlinear updates and a non-stationary policy, this algorithm adapts the Q-learning algorithm to optimize ESR for a potentially nonlinear welfare function.
Similarly, we run both RAVI and Welfare Q-learning on 20 random initializations of a $15\times 15$ scavenger hunt environment with six resources and 40% of the cells randomly populated with enemies, $T=20$ without reward discount ($\gamma=1$), and compare their scalarized total rewards at convergence. As shown in table 2, we again observe that RAVI is able to compute better performing policies across both nonlinear objectives compared to Welfare Q-learning.
Discretization of Accumulated Rewards.
We include an ablation study on the setting of the discretization interval length $\alpha$ in the appendix, in wihch we test values of $\alpha$ are substantially greater (thus making the algorithm more runtime efficient) than what would be necessary for worst-case theoretical guarantees from the previous analysis. Note that the empirical performance of RAVI is substantially better than is guaranteed by the previous theoretical analysis, for which substantially smaller $\alpha$ values would be necessary for worst-case guarantees.
7 Removing Knowledge of the Model
We have shown that RAVI can efficiently find an approximately optimal policy given access to a model of the environment. In this section we observe that the model can be jointly learned by extending the classical E3 algorithm [16] to the nonlinear multi-objective case by lifting the exploration algorithm to the multi-objective setting and using RAVI for the exploitation subroutine. We call the resulting combined algorithm Reward-Aware-Explore or Exploit (or RAEE for short). We briefly explain the high level ideas and state the main result here and defer further discussion to the appendix due to space constraints.
The algorithm consists of two stages, exploration and exploitation. Repeat the following stages, described here at a high level and in the appendix in more detail.
  • Explore. At a given state, choose the least experienced action and record the reward and transition. Continue in this fashion until reaching a known state, where we say a state is known if it has been visited sufficiently many times for us to have precise local statistics about the reward and transition functions at that state.
  • Exploit. Run RAVI from the current known state $s$ on the induced model comprising the known states and with a single absorbing state representing all unknown states. If the welfare obtained by this policy is within the desired error bound of $V^{*}(s,\mathbf{0},T)$ then we are done. Otherwise, compute a policy to reach an unknown state as rapidly as possible and resume exploring.
Theorem 2 (RAEE).
Let $V^{*}(s,0,T)$ denote the value function for the policy with the optimal expected welfare in the MOMDP $M$ starting at state $s$, with $\mathbf{0}\in\mathbb{R}^{d}$ accumulated reward and $T$ timesteps remaining. Then for a uniformly continuous welfare function $W$, there exists an algorithm $A$, taking inputs $\epsilon$, $\beta$, $\mathcal{S}$, $\mathcal{A}$, and $V^{*}(s,\mathbf{0},T)$, such that the total number of actions and computation time taken by $A$ is polynomial in $1/\epsilon$, $1/\beta$, $|\mathcal{S}|$, $|\mathcal{A}|$, the horizon time $T=1/(1-\gamma)$ and exponential in the number of objectives $d$, and with probability at least $1-\beta$, $A$ will halt in a state $s$, and output a policy $\hat{\pi}$, such that $V^{\hat{\pi}}_{M}(s,0,T)\geq V^{*}(s,0,T)-\epsilon$.
We provide additional details comparing our analysis with that of [16] in the appendix.
8 Conclusion and Future Work
Nonlinear prefernces in reinforcement learning are important, as they can encode fairness as the nonlinear balancing of priorities across multiple objectives or risk attitudes with respect to even a single objective. Stationary policies are not necessarily optimal for such objectives. We derived an extended form of Bellman optimality to characterize the structure of optimal policies conditional on accumulated reward. We introduced the RAVI and RAEE algorithms for efficiently computing an approximately optimal policy. Our primary contribution is a theoretical analysis, but we have also supplemented with an experimental demonstration.
Our results introduce several natural directions for future work. On the technical side, one could try to handle the case of stochastic reward functions or a large number of objectives. Another direction would involve incorporating function approximation with deep neural networks into the algorithms to enable scaling to even larger state spaces and generalizing between experiences. Our theory suggests that it may be possible to greatly enrich the space of possible policies that can be efficiently achieved in these settings by conditioning function approximators on accumulated reward rather than necessarily considering sequence models over arbitrary past trajectories; we see this as the most exciting next step for applications.
Appendix
Appendix A Proofs of Technical Lemmas
In this section, we provide proofs of the technical lemmas in Section 4 and Section 5.
Lemma 4.
Let $M$ be any MOMDP, and let $\pi$ be any policy in $M$. Assume the welfare function $W:\mathbb{R}^{d}\to\mathbb{R}$ is uniformly continuous and $\mathbf{R}(s,a)\in[0,1]^{d}~{}\forall s,a$. Then for all $\epsilon>0$, there exists $\delta_{\epsilon}>0$ such that if
$T\geq\left(\frac{1}{1-\gamma}\right)\log\left(\frac{1}{\delta_{\epsilon}(1-\gamma)}\right)$
then for any state $s,$
$V^{\pi}(s,\mathbf{0},T)\leq\lim_{T\to\infty}V^{\pi}(s,\mathbf{0},T)\leq V^{\pi}(s,\mathbf{0},T)+\epsilon.$
We call the value of the lower bound on $T$ given above the $\epsilon$-horizon time for the MOMDP $M$.
Proof.
[…]
$\bigg{\|}\sum_{k=1}^{\infty}\gamma^{k-1}\mathbf{R}(s_{k},a_{k})-\sum_{k=1}^{T}\gamma^{k-1}\mathbf{R}(s_{k},a_{k})\bigg{\|}$
$=$ $\bigg{\|}\sum_{k=T+1}^{\infty}\gamma^{k-1}\mathbf{R}(s_{k},a_{k})\bigg{\|}$
$=$ $\gamma^{T}\bigg{\|}\sum_{k=1}^{\infty}\gamma^{k-1}\mathbf{R}(s_{T+k},a_{T+k})\bigg{\|}$
$\leq$ $\gamma^{T}\left(\frac{1}{1-\gamma}\right)$
Solving
$\gamma^{T}\left(\frac{1}{1-\gamma}\right)\leq\delta_{\epsilon}$
for $T$ yields the desired bound on $T$. Since the inequality holds for every fixed trajectory, it also holds for the distribution over trajectories induced by any policy $\pi$. ∎
Lemma 1. Let $\mathcal{V}(s,\mathbf{R}(\tau),0)=W(\mathbf{R}(\tau))$ for all states $s$ and trajectories $\tau$. For every state $s$, history $\tau$, and $t>0$ time steps remaining, let
$\mathcal{V}(s,\mathbf{R}(\tau),t)=\max_{a}\mathbb{E}_{s^{\prime}}\left[\mathcal{V}(s^{\prime},\mathbf{R}(\tau)+\gamma^{T-t}\mathbf{R}(s,a),t-1)\right].$
Then $V^{*}(s,\tau,t)=\mathcal{V}(s,\mathbf{R}(\tau),t)$.
Proof.
We proceed by induction on $t$. In the base case of $t=0$, we have simply that $\mathcal{V}(s,\mathbf{R}(\tau),0)=W(\mathbf{R}(\tau))$ by the definition of $\mathcal{V}$. But when $t=0$, the trajectory has ended, so any $\tau^{\prime}$ in Definition 5 will be the empty trajectory $\tau_{0;0}$. So
$V^{*}(s,\tau,0)=W(\mathbf{R}(\tau)+\gamma^{T-t}\mathbf{R}(\tau_{0:0}))=W(\mathbf{R}(\tau)).$
Suppose the equality holds for all states $s$ and histories $\tau$ for up to $t-1$ time steps remaining, i.e.
$V^{*}(s,\tau,t-1)=\max_{a}\mathbb{E}_{s^{\prime}}\left[\mathcal{V}(s^{\prime},\mathbf{R}(\tau)+\gamma^{T-t}\mathbf{R}(s,a),t-2)\right].$
Then with $t$ steps remaining:
$V^{*}(s,\tau,t)=\max_{\pi}\sum_{\tau^{\prime}}Pr^{\pi}(\tau^{\prime})W(\mathbf{R}(\tau)+\gamma^{T-t}\mathbf{R}(\tau^{\prime}))$
where the sum is taken over all length-$t$ trajectories $\tau^{\prime}$ beginning at $s$, and $Pr^{\pi}(\tau^{\prime})$ is the probability that $\tau^{\prime}$ is traversed under policy $\pi$. Note that we can further decompose $\tau^{\prime}$ into $\tau^{\prime}=(s,a)\oplus\tau^{\prime\prime}$ where $(s,a)$ denotes the state-action pair corresponding to the current timestep, $\oplus$ concatenates, and $\tau^{\prime\prime}$ denotes a trajectory of length $t-1$ beginning at some state $s^{\prime}$ transitioned from state $s$ via taking action $a$. Then we can rewrite $V^{*}$ as
$V^{*}(s,\tau,t)=\max_{\pi}\sum_{a}\pi(a\mid s,\tau,t)~{}\mathbb{E}_{s^{\prime}}\bigg{[}\sum_{\tau^{\prime\prime}}Pr^{\pi}(\tau^{\prime\prime})W(\mathbf{R}(\tau)+\gamma^{T-t}\mathbf{R}(s,a)+\gamma^{T-t+1}\mathbf{R}(\tau^{\prime\prime}))\bigg{]}$
where the latter sum is taken over all length $t-1$ trajectories $\tau^{\prime\prime}$ beginning at $s^{\prime}$ and the expectation is taken over the environmental transitions to $s^{\prime}$ from $s$ given $a$.
Note that for all actions $a$, $\pi(a\mid s,\tau,t)\in[0,1]$ can be chosen independently of any terms that appears in $Pr^{\pi}(\tau^{\prime\prime})$ for any $\tau^{\prime\prime}$ by the decomposition of $\tau^{\prime}$ above. This implies that at the current timestep, the maximizer (or the optimal policy) should choose action $a$ that maximizes the quantity
$\mathbb{E}_{s^{\prime}}\bigg{[}\sum_{\tau^{\prime\prime}}Pr^{\pi}(\tau^{\prime\prime})W(\mathbf{R}(\tau)+\gamma^{T-t}\mathbf{R}(s,a)+\gamma^{T-t+1}\mathbf{R}(\tau^{\prime\prime}))\bigg{]}$
with probability 1. Therefore, the expression for $V^{*}(s,\tau,t)$ can be rewritten as
$V^{*}(s,\tau,t)$ $=\max_{a}\mathbb{E}_{s^{\prime}}\bigg{[}\max_{\pi}\sum_{\tau^{\prime\prime}}Pr^{\pi}(\tau^{\prime\prime})W(\mathbf{R}(\tau)+\gamma^{T-t}\mathbf{R}(s,a)+\gamma^{T-(t-1)}\mathbf{R}(\tau^{\prime\prime}))\bigg{]}$
$=\max_{a}\mathbb{E}_{s^{\prime}}\left[V^{*}(s^{\prime},\tau\oplus(s,a),t-1)\right]$
$=\max_{a}\mathbb{E}_{s^{\prime}}\left[\mathcal{V}(s^{\prime},\mathbf{R}(\tau)+\gamma^{T-t}\mathbf{R}(s,a),t-1)\right]$
$=\mathcal{V}(s,\mathbf{R}(\tau),t).$
where the second equality comes from the definition of $V^{*}$, the third equality comes from the inductive hypothesis, and the fourth equality comes from the definition of $\mathcal{V}$. ∎
Lemma 2. Let the welfare function $W:\mathbb{R}^{d}\to\mathbb{R}$ be uniformly continuous. Fix $s\in\mathcal{S}$ and $t\in\{0,1,\dots,T\}$, then for all $\epsilon>0$, there exists $\delta_{\epsilon}>0$ such that
$\bigg{|}V^{*}(s,\mathbf{R_{1}},t)-V^{*}(s,\mathbf{R_{2}},t)\bigg{|}<\epsilon\quad\text{if}\quad\|\mathbf{R_{1}}-\mathbf{R_{2}}\|<\delta_{\epsilon}.$
Proof.
Without loss of generality, assume $V^{*}(s,\mathbf{R_{1}},t)\geq V^{*}(s,\mathbf{R_{2}},t)$. Let $\pi_{1}$ be the induced optimal policy by $V^{*}(s,\mathbf{R_{1}},t)$, i.e.
$V^{*}(s,\mathbf{R_{1}},t)=V^{\pi_{1}}(s,\mathbf{R_{1}},t).$
Then by the optimality of $V^{*}$,
$V^{\pi_{1}}(s,\mathbf{R_{2}},t)$ $\leq V^{*}(s,\mathbf{R_{2}},t)$
$=\max_{\pi}\sum_{\tau}Pr^{\pi}[\tau]W(\mathbf{R_{2}}+\mathbf{R}(\tau)).$
Since $W$ is uniformly continuous, there exists $\delta_{\epsilon}>0$ such that $|W(\mathbf{x})-W(\mathbf{y})|<\epsilon$ if $\|\mathbf{x}-\mathbf{y}\|<\delta_{\epsilon}$. Then
$\bigg{|}V^{*}(s,\mathbf{R_{1}},t)-V^{*}(s,\mathbf{R_{2}},t)\bigg{|}$
$\leq$ $\bigg{|}V^{\pi_{1}}(s,\mathbf{R_{1}},t)-V^{\pi_{1}}(s,\mathbf{R_{2}},t)\bigg{|}$
$=$ $\bigg{|}\sum_{\tau}Pr^{\pi_{1}}[\tau]W(\mathbf{R_{1}}+\mathbf{R}(\tau))-\sum_{\tau}Pr^{\pi_{1}}[\tau]W(\mathbf{R_{2}}+\mathbf{R}(\tau))\bigg{|}$
$=$ $\sum_{\tau}Pr^{\pi_{1}}[\tau]\bigg{|}W(\mathbf{R_{1}}+\mathbf{R}(\tau))-W(\mathbf{R_{2}}+\mathbf{R}(\tau))\bigg{|}$
where the sum is over all $t$-trajectories $\tau$ that start in state $s$. If $\|\mathbf{R_{1}}-\mathbf{R_{2}}\|<\delta_{\epsilon}$ then we have
$\bigg{|}W(\mathbf{R_{1}}+\mathbf{R}(\tau))-W(\mathbf{R_{2}}+\mathbf{R}(\tau))\bigg{|}<\epsilon.$
Therefore,
$\bigg{|}V^{*}(s,\mathbf{R_{1}},t)-V^{*}(s,\mathbf{R_{2}},t)\bigg{|}<\epsilon$
since the sum of probabilities over all $t$-trajectories $\tau$ in $M$ that start in state $s$ induced by $\pi_{1}$ must add up to $1$. ∎
Lemma 3 (Approximation Error of RAVI)
For uniformly continuous welfare function $W$, for all $\epsilon>0$, there exists $\alpha_{\epsilon}$ such that
$V(s,\mathbf{R_{acc}},t)\geq V^{*}(s,\mathbf{R_{acc}},t)-t\epsilon$
$\forall s\in\mathcal{S},\mathbf{R_{acc}}\in\mathbb{R}^{d},t\in\{0,1,\dots,T\}$, where $V(s,\mathbf{R_{acc}},t)$ is computed by Algorithm 1.
Proof.
We prove this additive error bound by induction on the number of timesteps $t$ remaining in the task. Cleary $V(s,\mathbf{R_{acc}},0)=W(\mathbf{R_{acc}}),\forall s$ and $\mathbf{R_{acc}}$. Suppose that $V(s,\mathbf{R_{acc}},t)\geq V^{*}(s,\mathbf{R_{acc}},t)-t\epsilon$ for all $t=0,\dots,k-1$. Then consider with $k$ steps remaining.
$V(s,\mathbf{R_{acc}},k)=$ $\max_{a}\sum_{s^{\prime}}Pr(s^{\prime}|s,a)V(s^{\prime},f_{\alpha}(\mathbf{R_{acc}}+\gamma^{T-k}\mathbf{R}(s,a)),k-1)$
$\geq$ $\max_{a}\sum_{s^{\prime}}Pr(s^{\prime}|s,a)\times\bigg{(}V^{*}(s^{\prime},f_{\alpha}(\mathbf{R_{acc}}+\gamma^{T-k}\mathbf{R}(s,a)),k-1)-(k-1)\epsilon\bigg{)}.$
Since $W$ is uniformly continuous, by Lemma 2 there exists $\delta_{\epsilon}$ such that
$\bigg{|}V^{*}\bigg{(}s^{\prime},f_{\alpha}(\mathbf{R_{acc}}+\gamma^{T-k}\mathbf{R}(s,a)),k-1\bigg{)}-V^{*}\bigg{(}s^{\prime},\mathbf{R_{acc}}+\gamma^{T-k}\mathbf{R}(s,a),k-1\bigg{)}\bigg{|}<\epsilon$
if
$\bigg{\|}f_{\alpha}(\mathbf{R_{acc}}+\gamma^{T-k}\mathbf{R}(s,a))-(\mathbf{R_{acc}}+\gamma^{T-k}\mathbf{R}(s,a))\bigg{\|}<\delta_{\epsilon}.$
Hence, it suffices to choose $\alpha=\delta_{\epsilon}/d$, which implies
$\bigg{\|}f_{\alpha}(\mathbf{R_{acc}}+\gamma^{T-k}\mathbf{R}(s,a))-(\mathbf{R_{acc}}+\gamma^{T-k}\mathbf{R}(s,a))\bigg{\|}<(\delta_{\epsilon}/d)\cdot d=\delta_{\epsilon},$
and so
$V(s,\mathbf{R_{acc}},k)\geq$ $\max_{a}\sum_{s^{\prime}}Pr(s^{\prime}|s,a)V^{*}\bigg{(}s^{\prime},\mathbf{R_{acc}}+\gamma^{T-k}\mathbf{R}(s,a),k-1\bigg{)}-\epsilon-(k-1)\epsilon$
$=$ $V^{*}(s,\mathbf{R}_{acc},k)-k\epsilon.$
Appendix B Ablation Studies on Accumulated Rewards Discretization
We conduct further studies on the empirical effects of the discretization factor $\alpha$. For $W_{\text{SPF}}$, we set $\lambda=1$ (the smoothing factor). We gradually increase the disctretization interval of accumulated reward $\alpha$ gradually from $0.4$ to $1.2$ for a reward discount $\gamma=0.999$ for Taxi and $\gamma=0.99$ for Scavenger Hunt with the same settings as the experiments. As shown in Figure 4(a)(a) and Figure 4(b), we observe that the solution quality is often very high (better than baselines considered in the main body) even for discretization values that are substantially greater than those needed for theoretical guarantees (with $\epsilon=1$, we need $\alpha=2.5\times 10^{-3}$). Nonetheless the results are not necessarily monotone with respect to $\alpha$ in this regime.
(a) Nash Welfare in Taxi
(b) Cobb-Douglas in Scavenger Hunt
[…]
Figure 4: Ablation study on discretization factor $\alpha$.
This set of ablation studies demonstrates that the RAVI algorithm is capable of finding high-quality policies for optimizing expected scalarized return despite using a larger discretization parameter than that needed for our theoretical analysis. The practical takeaway is that, if one cannot afford the runtime for a very fine $\alpha$ with worst-case theoretical guarantees, one should instead try several much larger $\alpha$ values but still less than the smoothness of the objective. Due to the superlinear runtime scaling with respect to $\alpha$ for $d>1$, this can be substantially faster than a single run with much smaller $\alpha$.
Appendix C Details on the RAEE Algorithm
In this section, we give a detailed description of the RAEE algorithm, extending from [16] and demonstrating Theorem 2, the main theorem for the RAEE algorithm.
C.1 Introducing RAEE
The extension works as follows. The algorithm starts off by doing balanced wandering [16]: when encountering a new state, the algorithm selects a random action. However, when revisiting a previously visited state, it chooses the least attempted action from that state, resolving ties by random action selection. At each state-action pair $(s,a)$ it tries, the algorithm stores the reward $\mathbf{R}(s,a)$ received and an estimate of the transition probabilities $Pr(s^{\prime}\mid s,a)$ derived from the empirical distribution of next states reached during balanced wandering.
Next, we introduce the notion of a known state [16], which refers to a state that the algorithm has explored to the extent that the estimated transition probabilities for any action from that state closely approximate their actual values. Denote the number of times a state needs to be visited as $m_{known}$. We will specify the value later in our runtime characterization.
States are thus categorized into three groups: known states, which the algorithm has extensively visited and obtained reliable transition statistics; states that have been visited before but remain unknown due to limited trials and therefore unreliable data; and states that have not been explored at all. By the Pigeonhole Principle, accurate statistics will eventually accumulate in some states over time, leading to their becoming known. Let $S$ be the set of currently known states, the algorithm can build the current known-state MOMDP $M_{S}$ that is naturally induced on $S$ by the full MOMDP $M$ with all “unknown” states merged into a single absorbing state. Although the algorithm cannot access $M_{S}$ directly, it will have an approximation $\hat{M}_{S}$ by the definition of the known states. By the simulation lemma [16], $\hat{M}_{S}$ will be an accurate model in the sense that the expected $T$-step welfare of any policy in $\hat{M}_{S}$ is close to its expected $T$-step return in $M_{S}$. (Here $T$ is the horizon time.) Hence, at any timestep, $\hat{M}_{S}$ functions as an incomplete representation of $M$, specifically focusing on the part of $M$ that the algorithm possesses a strong understanding of.
This is where we insert RAVI. The algorithm performs the two off-line optimal policy computations; i) first on $\hat{M}_{S}$ using RAVI to compute an exploitation policy that yields an approximately optimal welfare and ii) second performing traditional value iteration on $\hat{M}^{\prime}_{S}$, which has the same transition probabilities as $\hat{M}_{S}$, but different payoffs: in $\hat{M}^{\prime}_{S}$, the absorbing state (representing “unknown” states) has scalar reward $1$ and all other states have scalar reward 00. The optimal policy in $\hat{M}^{\prime}_{S}$ simply exits the known model as rapidly as possible, rewarding exploration.
By the explore or exploit lemma [16], the algorithm is guaranteed to either output a policy with approximately optimal return in $M$, or to improve the statistics at an unknown state. Again by the Pigeonhole Principle, a new state becomes known after the latter case occurs for some finite number of times, and thus the algorithm is always making progress. In the worst case, the algorithm builds a model of the entire MOMDP $M$. Having described the elements, we now outline the entire extended algorithm, where the notations are consistent with [16].
RAEE Algorithm:
  • (Initialization) Initially, the set $S$ of known states is empty.
  • (Balanced Wandering) Any time the current state is not in $S$, the algorithm performs balanced wandering
  • (Discovery of New Known States) Any time a state $s$ has been visited $m_{known}$ [16] times during balanced wandering, it enters the known set $S$, and no longer participates in balanced wandering.
  • (Off-line Optimizations) Upon reaching a known state $s\in S$ during balanced wandering, the algorithm performs the two off-line optimal policy computations on $\hat{M}_{S}$ and $\hat{M}^{\prime}_{S}$ described above:
    • (Attempted Exploitation) Use RAVI algorithm to compute an $\epsilon/2$-optimal policy on $\hat{M}_{S}$. If the resulting exploitation policy $\hat{\pi}$ achieves return from $s$ in $\hat{M}_{S}$ that is at least $V^{*}(s,\mathbf{0},T)-\epsilon/2$, the algorithm halts and outputs $\hat{\pi}$.
    • (Attempted Exploration) Otherwise, the algorithm executes the resulting exploration policy derived from the off-line computation on $\hat{M}^{\prime}_{S}$ for $T$ steps in $M$.
  • (Balanced Wandering) Any time an attempted exploitation or attempted exploration visits a state not in $S$, the algorithm resumes balanced wandering.
This concludes the description of the algorithm.
C.2 Runtime Analysis
In this subsection, we comment on additional details of the analysis of Theorem 2, the main theorem for the RAEE algorithm.
Theorem 2 (RAEE). Let $V^{*}(s,0,T)$ denote the value function for the policy with the optimal expected welfare in the MOMDP $M$ starting at state $s$, with $\mathbf{0}\in\mathbb{R}^{d}$ accumulated reward and $T$ timesteps remaining. Then for a uniformly continuous welfare function $W$, there exists an algorithm $A$, taking inputs $\epsilon$, $\beta$, $|\mathcal{S}|$, $|\mathcal{A}|$, and $V^{*}(s,\mathbf{0},T)$, such that the total number of actions and computation time taken by $A$ is polynomial in $1/\epsilon$, $1/\beta$, $|\mathcal{S}|$, $|\mathcal{A}|$, the horizon time $T=1/(1-\gamma)$ and exponential in the number of objectives $d$, and with probability at least $1-\beta$, $A$ will halt in a state $s$, and output a policy $\hat{\pi}$, such that $V^{\hat{\pi}}_{M}(s,0,T)\geq V^{*}(s,0,T)-\epsilon$.
We begin by defining approximation of MOMDPs. Think of $M$ as the true MOMDP, that is, a perfect model of transition probabilities. Think of $M^{\prime}$, on the other hand, as the current best estimate of an MOMDP obtained through exploration. In particular, $M^{\prime}$ in the RAEE algorithm will consist of the set of known states.
Definition 8.
[16] Let $M$ and $M^{\prime}$ be two MOMDPs over the same state space $\mathcal{S}$ with the same deterministic reward function $\mathbf{R}(s,a)$. $M^{\prime}$ is an $\alpha-$approximation of $M$ if for any state $s$ and $s^{\prime}$ and any action $a$,
$Pr_{M}(s^{\prime}\mid s,a)-\alpha\leq Pr_{M^{\prime}}(s^{\prime}\mid s,a)\leq Pr_{M}(s^{\prime}\mid s,a)+\alpha,$
where the subscript denotes the model.
We now extend the Simulation Lemma [16], which tells us how close the approximation of an MOMDP needs to be in order for the expected welfare, or ESR, of a policy to be close in an estimated model. The argument is similar to [16], but in our case the policy may not be stationary and we want to bound the deviation in expected welfare rather than just accumulated reward.
Recall that $G^{T}_{max}$ is defined to be the maximum possible welfare achieved on a $T$-step trajectory - $G^{T}_{max}$ is at most $W(\mathbf{T})$ where $\mathbf{T}$ equals $T$ times the identity vector, in our model.
Lemma 5 (Extended Simulation Lemma).
Let $M^{\prime}$ be an $O((\epsilon/|\mathcal{S}||\mathcal{A}|TG^{T}_{\max})^{2})$-approximation of $M$. Then for any policy $\pi$, any state $s$, and horizon time $T$, we have
$V^{\pi}_{M}(s,\tau_{0;0},T)-\epsilon\leq V^{\pi}_{M^{\prime}}(s,\tau_{0;0},T)\leq V^{\pi}_{M}(s,\tau_{0;0},T)+\epsilon.$
Proof.
Fix a policy $\pi$ and a start state $s$. Let $M^{\prime}$ be an $\alpha$-approximation of $M$ (we will later show that $\alpha$ has the same bound as the Lemma statement). Call the transition probability from a state $s$ to a state $s^{\prime}$ under action $a$ to be $\beta$-small in $M$ if $Pr_{M}(s^{\prime}\mid s,a)\leq\beta$. Then the probability that a $T$-trajectory $\tau$ starting from a state $s$ following policy $\pi$ contains at least one $\beta$-small transition is at most $\beta|\mathcal{S}||\mathcal{A}|T$. This is because the total probability of all $\beta$-small transitions in $M$ is at most $\beta|\mathcal{S}||\mathcal{A}|$ (assuming all transition probabilities are $\beta$-small), and there are $T$ timesteps. Note that in our case, the optimal policy may not be necessarily stationary, thus the agent does not necessarily choose the same action (and hence the same transition probability) upon revisiting any state. So we cannot bound the total probability by $\beta|\mathcal{S}|$ like in the original proof.
The total expected welfare of the trajectories of $\pi$ that consist of at least one $\beta$-small transition of $M$ is at most $\beta|\mathcal{S}||\mathcal{A}|TG^{T}_{\max}$. Recall that $M^{\prime}$ be an $\alpha$-approximation of $M$. Then for any $\beta$-small transition in $M$, we have $Pr_{M^{\prime}}(s^{\prime}\mid s,a)\leq\alpha+\beta$. So the total welfare of the trajectories of $\pi$ that consist of at least one $\beta$-small transition of $M$ is at most $(\alpha+\beta)|\mathcal{S}||\mathcal{A}|TG^{T}_{\max}$. We can thus bound the difference between $V^{\pi}_{M}(s,\tau_{0;0},T)$ and $V^{\pi}_{M^{\prime}}(s,\tau_{0;0},T)$ restricted to these trajetories by $(\alpha+2\beta)|\mathcal{S}||\mathcal{A}|TG^{T}_{\max}$. We will later choose $\beta$ and bound this value by $\epsilon/4$ to solve for $\alpha$.
Next, consider trajectories of length $T$ starting at $s$ that do not contain any $\beta$-small transitions, i.e. $Pr_{M}(s^{\prime}\mid s,a)>\beta$ in these trajectories. Choose $\Delta=\alpha/\beta$, we may write
$(1-\Delta)Pr_{M}(s^{\prime}\mid s,a)$
$\qquad\leq Pr_{M^{\prime}}(s^{\prime}\mid s,a)$
$\qquad\leq(1+\Delta)Pr_{M}(s^{\prime}\mid s,a).$
because $M^{\prime}$ is an $\alpha$-approximation of $M$ and $\beta\leq 1$. Thus for any $T$-trajectory $\tau$ that does not cross any $\beta$-small transitions under $\pi$, we have
$(1-\Delta)^{T}Pr_{M}^{\pi}[\tau]\leq Pr_{M^{\prime}}^{\pi}[\tau]\leq(1+\Delta)^{T}Pr_{M}^{\pi}[\tau],$
which follows from the definition of the probability along a trajectory and the fact that $\pi$ is the same policy in all terms. Since we assume reward functions are deterministic in our case, for any particular $T$-trajectory $\tau$, we also have
$W_{M}(\mathbf{R}(\tau))=W_{M^{\prime}}(\mathbf{R}(\tau)).$
Since these hold for any fixed $T$-trajectory that does not traverse any $\beta$-small transitions in $M$ under $\pi$, they also hold when we take expectations over the distributions over such $T$-trajectories in $M$ and $M^{\prime}$ induced by $\pi$. Thus
$(1-\Delta)^{T}V^{\pi}_{M}(s,\tau_{0;0},T)-\frac{\epsilon}{4}$
$\qquad\leq V^{\pi}_{M^{\prime}}(s,\tau_{0;0},T)$
$\qquad\leq(1+\Delta)^{T}V^{\pi}_{M}(s,\tau_{0;0},T)+\frac{\epsilon}{4}.$
where the $\epsilon/4$ terms account for the contributions of the $T$trajectories that traverse at least one $\beta$-small transitions under $\pi$. It remains to show how to choose $\alpha$, $\beta$, and $\Delta$ to obtain the desired approximation. For the upper bound, we solve for
$(1+\Delta)^{T}V^{\pi}_{M}(s,\tau_{0;0},T)\leq V^{\pi}_{M}(s,\tau_{0;0},T)+\frac{\epsilon}{4}$
$\implies(1+\Delta)^{T}\leq 1+\epsilon/(4G^{T}_{\max}).$
By taking $\log$ on both sides and using Taylor expansion, we can upper bound $\Delta$.
$T\Delta/2\leq\epsilon/(4G^{T}_{\max})$
$\qquad\implies\Delta\leq\epsilon/(2TG^{T}_{\max}).$
Choose $\beta=\sqrt{\alpha}$. Then
$\begin{cases}(\alpha+2\beta)|\mathcal{S}||\mathcal{A}|TG^{T}_{\max}\leq 3\sqrt{\alpha}|\mathcal{S}||\mathcal{A}|TG^{T}_{\max}\leq\epsilon/4\\ \Delta=\sqrt{\alpha}\leq\epsilon/(2TG^{T}_{\max})\end{cases}$
Choosing $\alpha=O((\epsilon/|\mathcal{S}||\mathcal{A}|TG^{T}_{\max})^{2})$ solves the system. The lower bound can be handled similarly, which completes the proof of the lemma. ∎
We now define a “known state.” This is a state that has been visited enough times, and its actions have been trialed sufficiently many times, that we have accurate estimates of the transition probabilities from this state.
Definition 9.
[16] Let $M$ be an MOMDP. A state $s$ of $M$ is considered known if it has be visited a number of times equal to
$m_{known}=O((|\mathcal{S}||\mathcal{A}|TG^{T}_{\max}/\epsilon)^{4}|\mathcal{A}|\log(1/\delta)).$
By applying Chernoff bounds, we can show that if a state has been visited $m_{known}$ times then its empircal estimation of the transition probabilities satisfies the accuracy required by the Lemma 5.
Lemma 6.
[16] Let $M$ be an MOMDP. Let $s$ be a state of $M$ that has been visited at least $m$ times, with each action having been executed at least $\lfloor m/|\mathcal{A}|\rfloor$ times. Let $\hat{Pr}(s^{\prime}\mid s,a)$ denote the empirical transition probability estimates obtained from the $m$ visits to $s$. Then if
$m=O((|\mathcal{S}||\mathcal{A}|TG^{T}_{\max}/\epsilon)^{4}|\mathcal{A}|\log(1/\delta)),$
then with probability $1-\delta$, we have
$|\hat{Pr}(s^{\prime}\mid s,a)-Pr(s^{\prime}\mid s,a)|=O((\epsilon/|\mathcal{S}||\mathcal{A}|TG^{T}_{\max})^{2})$
for all $s^{\prime}\in\mathcal{S}$.
Proof.
The sampling version of Chernoff bounds states that if the number of independent, uniformly random samples $n$ that we use to estimate the fraction of a population with certain property $p$ satisfies
$n=O\left(\frac{1}{\alpha^{2}}\log\left(\frac{1}{\delta}\right)\right),$
the our estimate $\bar{X}$ satisfies
$\bar{X}\in[p-\alpha,p+\alpha]~{}\text{with probability $1-\delta$}.$
By the Extended Simulation Lemma, it suffices to choose $\alpha=O((\epsilon/|\mathcal{S}||\mathcal{A}|TG^{T}_{\max})^{2})$.
Note that we need to insert an extra factor of $|\mathcal{A}|$ compared to the original analysis since we treat the size of the action space as a variable instead of a constant, and a state is categorized as “known” only if the estimates of transition probability of all actions are close enough. ∎
We have specified the degree of approximation required for sufficient simulation accuracy. It remains to directly apply the Explore or Exploit Lemma from [16] to conclude the analysis.
Lemma 7 (Explore or Exploit Lemma [16]).
Let $M$ be any MOMDP, let $S$ be any subset of the states of $M$, and let $M_{S}$ be the MOMDP on $M$. For any $s\in\mathcal{S}$, any $T$, and any $01>\alpha>$ 0, either there exists a policy $\pi$ in $M_{S}$ such that $V^{\pi}_{M_{S}}(s,\tau_{0;0},T)\geq V^{*}_{M}(s,\mathbf{0},T)-\alpha$, or there exists a policy $\pi$ in $M_{S}$ such that the probability that a $T$-trajectory following $\pi$ will lead to the exit state exceeds $\alpha/G^{T}_{\max}$.
This lemma guarantees that either the $T$step return of the optimal exploitation policy in the simulated model is very close to the optimal achievable in $M$, or the agent choosing the exploration policy can reach a previously unknown state with significant probability.
Note that unlike the original E3 algorithm, which uses standard value iteration to compute the exactly optimal policy (optimizing a linear function of a scalar reward) on the sub-model $\hat{M}_{S}$, we use our RAVI algorithm to find an approximately optimal policy. Therefore, we need to allocate $\epsilon/2$ error to both the simulation stage and the exploitation stage, which gives a total of $\epsilon$ error for the entire algorithm.
It remains to handle the failure parameter $\beta$ in the statement of the main theorem, which can be done similarly to [16]. There are two sources of failure for the algorithm:
  • The algorithm’s estimation of the next state distribution for some action at a known state is inaccurate, resulting $\hat{M}_{S}$ being an inaccurate model of $M_{S}$.
  • Despite doing attempted explorations repeatedly, the algorithm fails to turn a previously unknown state into a known state because of an insufficient number of balanced wandering.
It suffices to allocate $\beta/2$ probabilities to each source of failure. The first source of failure are bounded by Lemma 6. By choosing $\beta^{\prime}=\beta/(2|\mathcal{S}|)$, we ensure that the probability that the first source of failure happens to each of the known states in $M_{S}$ is sufficiently small such that the total failure probability is bounded by $\beta/2$ for $M_{S}$.
For the second source of failure, by the Explore or Exploit Lemma, each attempted exploration results in at least one step of balanced wandering with probability at least $\epsilon/(2G^{T}_{\max})$, i.e. when this leads the agent to an unknown state. The agent does at most $|\mathcal{S}|m_{known}$ steps of balanced wandering, since this makes every state known. By Chernoff bound, the probability that the agent does fewer than $|\mathcal{S}|m_{known}$ steps of balanced wandering (attempted explorations that actually leads to an unknown state) will be smaller than $\beta/2$ if the number of attempted explorations is
$O((G^{T}_{\max}/\epsilon)|\mathcal{S}|\log(|\mathcal{S}|/\beta)m_{known}),$
where $m_{known}=O(((|\mathcal{S}||\mathcal{A}|TG^{T}_{\max})/\epsilon)^{4}|\mathcal{A}|\log(|\mathcal{S}|/\beta))$ (recall we choose $\beta^{\prime}=\beta/(2|\mathcal{S}|)$ for the first source of failure).
Thus, the total computation time is bounded by $O\left(|\mathcal{S}|^{2}|\mathcal{A}|(T/\alpha_{T\epsilon/2})^{d}\right)$ (the time required for RAVI during off-line computations with $\epsilon/2$ precision by Lemma 3) times $T$ times the maximum number of attempted explorations, giving
$O((|\mathcal{S}|^{3}|\mathcal{A}|TG^{T}_{\max}/\epsilon)(T/\alpha_{T\epsilon/2})^{d}\log(|\mathcal{S}|/\beta)m_{known}).$
This concludes the proof of Theorem 2.